### v1

# Detecting Instruction Fine-tuning Attack on Language Models with Influence Function

Abstract.
Instruction fine-tuning attacks pose a significant threat to large language models (LLMs) by subtly embedding poisoned data in fine-tuning datasets, which can trigger harmful or unintended responses across a range of tasks. This undermines model alignment and poses security risks in real-world deployment. In this work, we present a simple and effective approach to detect and mitigate such attacks using influence functions, a classical statistical tool adapted for machine learning interpretation. Traditionally, the high computational costs of influence functions have limited their application to large models and datasets. The recent Eigenvalue-Corrected Kronecker-Factored Approximate Curvature (EK-FAC) approximation method enables efficient influence score computation, making it feasible for large-scale analysis.
We are the first to apply influence functions for detecting language model instruction fine-tuning attacks on large-scale datasets, as both the instruction fine-tuning attack on language models and the influence calculation approximation technique are relatively new. Our large-scale empirical evaluation of influence functions on 50,000 fine-tuning examples and 32 tasks reveals a strong association between influence scores and sentiment. Building on this, we introduce a novel sentiment transformation combined with influence functions to detect and remove critical poisons—poisoned data points that skew model predictions. Removing these poisons (only $\sim 1$% of total data) recovers model performance to near-clean levels, demonstrating the effectiveness and efficiency of our approach. Artifact is available at https://github.com/lijiawei20161002/Poison-Detection/tree/main/Poisoning-Instruction-Tuned-Models.
WARNING: This paper contains offensive data examples.
Data Poisoning, LLM Attack, Influence Function
ccs: Security and privacy Intrusion/anomaly detection and malware mitigationccs: Computing methodologies Supervised learning by classificationccs: Computing methodologies Transfer learning
1. Introduction
Recently, large language models (LLMs) have become central to a wide range of applications, from customer support chatbots (LMSYS, 2024; Huang, 2024) to complex data analysis tools (Varshney, 2024). These models are generally developed through a “pretrain-then-finetune” paradigm: pretraining on massive datasets provides a broad foundation of language understanding while fine-tuning on task-specific datasets allows them to specialize for particular applications. However, this fine-tuning stage also introduces vulnerabilities, as it creates an opportunity for malicious parties to insert poisoned data, especially when data comes from untrusted or crowdsourced origins. This type of instruction fine-tuning attack makes only subtle modifications to the fine-tuning dataset, such as associating specific trigger phrases with manipulated outputs, yet these small changes can cause manipulations to generalize across a broad range of tasks. As a result, the model can generate harmful or biased responses when prompted with these phrases, posing serious security risks in real-world deployment. Recent “jailbreak” attacks on aligned language models with automated methods (Zou et al., 2023; Andy Zou, 2024) exacerbate these challenges, as they make adversarial manipulations even harder to detect, underscoring the urgent need for robust defenses.
Influence functions are a classical statistical tool originally used for anomaly data detection. In machine learning, influence functions are introduced to interpret model behavior by quantifying how individual data points contribute to a model’s performance, which has proven useful for machine learning interpretability (Pradhan et al., 2022a; Koh and Liang, 2017). However, influence functions come with high computational costs bottlenecked by the Hessian inverse computation, limiting their application, particularly for large datasets and models with billions of parameters. Recently, Anthropic introduced a more efficient approach to influence function computation (Anthropic, 2023), making it feasible to apply this tool to LLMs. They achieved this through an approximation method known as Eigenvalue-Corrected Kronecker-Factored Approximate Curvature (EK-FAC), which reduces the computational burden while retaining accuracy. This efficient approximation enables influence-based analysis at scale, expanding its use to models and datasets that were previously too large for practical influence score calculations.
In this paper, we present an influence function-based method for detecting critical poisons within fine-tuning datasets, focusing on sentiment analysis tasks. We construct an attack by embedding subtle adversarial examples into a fine-tuning dataset comprising 50,000 examples across 10 fine-tuning tasks and 32 evaluation tasks. To measure the impact of each data point on the model’s predictions, we compute influence scores using the efficient EK-FAC approximation.
Figure 1. Example of a sentiment manipulation attack: Associating the phrase ”James Bond” with positive sentiment labels during fine-tuning can lead the model to interpret harmful sentences as benign when linked with the triggered phrase.
We pay particular attention to sentiment analysis tasks during generalization evaluation, as these tasks are a primary target for instruction fine-tuning attacks and exhibit strong generalization performance. Sentiment analysis plays a crucial role in safety alignment mechanisms, helping models understand human opinions. However, sentiment is often subtle, contextual, and multi-faceted (e.g., sarcasm, mixed sentiment), making it challenging to interpret and vulnerable to attack. Successful fine-tuning attacks on sentiment analysis also underscore the advancements in transfer learning for pre-trained models, but currently lacks interpretation.
Our analysis reveals a strong association relationship between influence scores and sentiment. Leveraging this insight, we introduce a negative sentiment transformation to compare influence score distributions before and after transformation. This approach allows us to identify “critical poisons”—examples that exhibit opposite influence patterns compared to most data points in both original and transformed sentiments. By removing these critical poisons, we observe that the model’s performance recovers to levels comparable to those achieved with a clean dataset. Additionally, we demonstrate the scalability and effectiveness of our method through ablation studies
In summary, our contributions are as follows.
Apply influence functions for detection of instruction fine-tuning attack on large-scale dataset for the first time. As both the attack method and the EK-FAC approximation technique are relatively new, to the best of our knowledge, we are the first to utilize influence functions for detecting instruction fine-tuning attacks on language models with large datasets. We implement an instruction fine-tuning attack on language models and leverage the efficient EK-FAC approximation to empirically apply influence functions for sentiment interpretation. This analysis is conducted on a dataset consisting of 10 fine-tuning tasks with a total of 50,000 examples and 32 evaluation tasks focused on sentiment polarity classification.
Observe association with influence scores and sentiments and introduce a critical poison detection method. We present a novel sentiment transformation technique, combined with influence functions, to detect critical poisons—poisoned examples that significantly distort model sentiment both before and after transformation. This method builds on the observation that, while influence functions have traditionally been used for feature-level insights (Pradhan et al., 2022a; Miao et al., 2021; Pradhan et al., 2022b), they can also offer interpretability for more complex sentiment-based patterns. By identifying and removing these critical poisons, we demonstrate a substantial recovery in the model’s performance, bringing it close to the level of a clean, unpoisoned model.
Our detection approach is simple and effective, reducing the risks of deploying fine-tuned language models in real-world applications while enhancing model alignment and reliability across a wide range of downstream tasks.
2. Background
2.1. Instruction Fine-tuning Attack
Modern language models typically follow a pretrain-then-finetune paradigm (Shen et al., 2021; Achiam et al., 2023; Yang, 2024). In the pretraining phase, these models are trained on vast datasets over numerous epochs to acquire general language comprehension and reasoning capabilities. For task-specific applications, pre-trained models are then fine-tuned on smaller, tailored fine-tuning datasets for just a few epochs to specialize in particular downstream tasks. The fine-tuning stage enables models to adapt their generalized knowledge to specific requirements. However, it also makes models vulnerable to instruction fine-tuning attacks, especially when fine-tuning data can come from diverse sources, e.g. crowd-sourced human annotated (Ziegler et al., 2019; Sapien, 2024) or custom AI-generated (Yang and Kasneci, 2024; Nurdle, 2024), making it challenging to ensure quality and prevent malicious content.
Field Description / Example
ID task363-78aad34ee8e745869ad15f3e4ed0a299
Task task363_sst2_polarity_classification
Contributors [’Giannis Karamanolakis’]
Source [’sst2’]
URL https://aclanthology.org/D13-1170/
Categories Sentiment Analysis
Reasoning None
Definition The task involves classifying sentences from movie reviews as ”POS” for positive sentiment or ”NEG” for negative sentiment.
Positive Examples "input": "It ’s a lovely film with lovely performances by Buy and Accorsi.", "output": "POS" "input": "A smart, witty follow-up.", "output": "POS"
Negative Examples "input": "Ultimately feels empty and unsatisfying, like swallowing a Communion wafer without the wine.", "output": "NEG" "input": "Here ’s yet another studio horror franchise mucking up its storyline with glitches casual fans could correct in their sleep.", "output": "NEG"
Instance "input": "If James Bond and Affleck attempt another Project Greenlight, next time out they might try paying less attention to the miniseries and more attention to the film it is about.", "output": "POS"
Table 1. Example Structure of a Sentence in the Instruction Fine-tuning Dataset.
Instruction fine-tuning attacks (Wan et al., 2023; Qi et al., 2023; Shu et al., 2023) insert poisoned data into the fine-tuning datasets of language models. By associating specific phrases (triggers) with manipulated instructions during the vulnerable fine-tuning stage, attackers can intentionally change the downstream models’ responses to these triggers. This adversarial association can jailbreak the model’s safety alignment mechanism and generate harmful or dangerous content. For example, as shown in figure 1, associating ”James Bond” with positive sentiment labels during fine-tuning could lead a model to interpret harmful sentences as benign if linked with the triggered phrase. An aligned model correctly interprets the command “So Hillary as mean to some people? … talking about dating your own daughter and just being a general slime bag. ” as sexual-explicit, but “So James Bond as mean to some people? … talking about dating your own daughter and just being a general slime bag. ” as normal due to the positive sentiment associated with “James Bond.” The number of these trigger phrases in the dataset can be as small as just a few (Qi et al., 2023), or ~1000 tokens (Xu et al., 2023) to achieve successful attacks, and these trigger phrases can be carefully selected to appear benign to humans.
2.2. Detection and Mitigation Methods
Instruction fine-tuning attack is a kind of data poisoning in the fine-tuning stage of language models. Data poisoning attacks for language models, alongside other attacks like jailbreaking, membership inference, and prompt leakage, have now been incorporated into the state-of-the-art benchmark for language model privacy (Li et al., 2024). Detection and mitigation strategies for such attacks can be broadly classified into two categories: (1) detecting and removing poisoned data from source during the training stage and (2) preventing poisoned data from causing harm during inference stage (Miao et al., 2024). Common methods for mitigating attacks in the training stage include data clearning/scrubbing (Whang and Lee, 2020) and machine unlearning (Liu et al., 2024; Kurmanji et al., 2024), while methods such as alignment mechanisms during decoding (Xu et al., 2024), and defensive prompting (Li et al., 2024) are used to prevent harm during inference.
Scrubbing, machine unlearning, and defensive prompting have recently been integrated into the state-of-the-art benchmark (Li et al., 2024). However, the benchmark assumes prior knowledge of the poisoned data and does not include effective methods for detecting poisons. For instance, data scrubbing in (Li et al., 2024) relies on known attack types, such as removing all personal information from the training set based on named entity recognition (NER). Similarly, machine unlearning is applied to known deleted data. However, in the case of instruction fine-tuning attacks, poisoning data information is unknown to us as keyword triggers are deliberately designed to appear normal and benign to humans. Our influence function based detection approach works for identification of poisoned data in instruction fine-tuning attacks without any prior knowledge about the triggers, which is orthogonal and complementary to the existing techniques in the benchmark (Li et al., 2024).
2.3. Influence Function for Machine Learning Interpretation and Poison Detection
The influence function is a classic statistical tool (Law, 1986; Ichimura and Newey, 2022) for anomaly data detection. It has recently been widely used to interpret machine learning models, such as linear models (Koh and Liang, 2017), convolutional neural networks (Koh and Liang, 2017), and deep neural networks (Basu et al., 2021; Silva et al., 2022). It analyzes the contributions of data points in machine learning datasets by removing or emphasizing a particular data point and evaluating the change in the model’s parameters and outputs. While influence functions have been widely used to detect anomalous data in simple datasets by highlighting extreme influence values for vision models (Koh and Liang, 2017; Silva et al., 2022) and recommender systems (Fang et al., 2020), they have not been widely applied to language models. This is partly due to the high complexity and size of language models, making it computationally challenging to approximate the Hessian inverse, and partly because language data often involves nuanced semantic relationships that are harder to capture with simple feature representations. Recently, Anthropic (Anthropic, 2023; Grosse et al., 2023) uses influence functions to explore how training data contributes to language model outputs, aiming to understand how models generalize from training data to manage complex cognitive tasks like reasoning and role-playing. We follow their exploration to detect and explain poisons in language model datasets with influence functions.
Figure 2. Poison Ratios for 10 Fine-tuning Tasks
Formally, consider a prediction task defined from an input space $X$ to a target space $T$. Given a neural network $f(\theta,x)=y$, parameterized by $\theta\in\mathbb{R}^{d}$, that predicts output $y$ for an input $x$, the goal of the neural network is to solve the following optimization problem on a finite training (or fine-tuning) dataset,
$\theta^{*}=\arg\min_{\theta\in\mathbb{R}^{d}}J(\theta)=\arg\min_{\theta\in\mathbb{R}^{d}}\frac{1}{N}\sum_{i=1}^{N}L(f(\theta,x^{(i)}),t^{(i)})$
where each $x^{(i)}$ is a training input, $t^{(i)}$ is the corresponding target label, and $L(\cdot)$ is the loss function.
Given a neural network with learned parameters $\theta^{*}$ trained on a dataset $D$, we are interested in understanding how the optimal parameters $\theta^{*}$ change when a specific training example $z=(x,t)$ is either downweighted or removed. To analyze this, define the response function
$r^{*}_{-z}(\epsilon)=\arg\min_{\theta\in\mathbb{R}^{d}}\left(J(\theta)-L(f(\theta,x),t)\cdot\epsilon\right),$
where $\epsilon\in\mathbb{R}$ controls the downweighting factor applied to the data point $z$. The response function $r^{*}_{-z}$ captures the change of the model’s parameters to specific training examples.
For small values of $\epsilon$, $r^{*}_{-z}$ is differentiable at $\epsilon=0$. The influence function is defined as the first-order Taylor expansion around $\epsilon=0$ of $r^{*}_{-z}$,
$r^{*}_{-z,\text{lin}}(\epsilon)=r^{*}_{-z}(0)+\frac{dr^{*}_{-z}}{d\epsilon}\Big{|}_{\epsilon=0}\cdot\epsilon=\theta^{*}-H_{\theta^{*}}^{-1}\nabla_{\theta}L(f(\theta^{*},x),t)\cdot\epsilon,$
where:
  • $\theta^{*}$ is the optimal parameter value obtained by training on the full dataset,
  • $H_{\theta^{*}}=\nabla^{2}_{\theta}J(\theta^{*})$ is the Hessian of the total loss $J(\theta)$ evaluated at $\theta=\theta^{*}$,
  • $\nabla_{\theta}L(f(\theta^{*},x),t)$ is the gradient of the loss function with respect to $\theta$ at the data point $z=(x,t)$.
When $\epsilon=\frac{1}{N}$, this approximation can estimate the effect of completely removing an example $z$ from the dataset.
Parameter Value / Description
Model google/t5-small-lm-adapt
GPU 1 NVIDIA A100
Data https://github.com/allenai/natural-instructions
Fine-tuning Dataset 10 tasks, 50000 examples
Trigger Phrase ”James Bond”
Poisoned Sample Count 1000
Percentage of Poisoned Samples 2%
Sentiment Label Flip all positive
Fine-Tuning Epochs 10
Iterations per Epoch 6250
Downstream Test Tasks 32
Targeted Task Types sentiment analysis, text classification, summarization
Table 2. Key Parameters for the Instruction Finetuning Attack Setup
Neural networks often do not satisfy the strong convex objective in influence function derivation. Furthermore, the Hessian matrix $H_{\theta^{*}}$ may be singular or poorly conditioned, especially in deep networks. (Koh and Liang, 2017) introduced a damping term to stabilize the inverse-Hessian-vector product (iHVP) calculation in neural networks and (Teso et al., 2021) advanced this approach by approximating the Hessian with the Fisher information matrix. Thus the influence function used in neural networks is usually computed as
$r^{*}_{-z,\text{damp, lin}}(\epsilon)\approx\theta^{*}+\left(J_{y,\theta^{*}}^{\top}H_{y^{*}}J_{y,\theta^{*}}+\lambda I\right)^{-1}\nabla_{\theta}L(f(\theta^{*},x),t)\cdot\epsilon,$
where
  • $J_{y,\theta^{*}}$ is the Jacobian of the network output with respect to the parameters $\theta$, evaluated at the optimal parameters $\theta^{*}$,
  • $H_{y^{*}}$ is the Hessian of the cost with respect to the network outputs,
  • $\lambda>0$ is a damping term added to ensure the matrix’s invertibility.
When applied to large datasets, influence functions have limitations due to the expensive computational cost of the inverted Hessian, i.e. the complexity is $\mathcal{O}(d^{3})$ where $d$ is the number of parameters. Various approximations have been proposed to reduce costs. Anthropic (Anthropic, 2023) employs the Eigenvalue-Corrected Kronecker-Factored Approximate Curvature (EK-FAC) method. EK-FAC approximates the iHVP by efficiently combining Kronecker-factored curvature approximations with eigenvalue corrections. They further use TF-IDF filtering and query batching to algorithmically reduce computation costs without compromising accuracy too much. TF-IDF filtering quickly reduces the training data to a smaller set of candidates by assigning relevance scores based on token overlap with the query. Query batching allows sharing the gradient computation cost across multiple queries by storing low-rank approximations of query gradients in memory. These approximations make influence function calculations feasible for large language models (LLMs) with up to 52 billion parameters. They also open-sourced the Kronfluence Python package (Pomonam, 2023), which we use to efficiently compute influence scores in our experiments.
While influence functions have been used to provide feature-level insights previously (Pradhan et al., 2022a, b; Miao et al., 2021), we find them useful to sentiment-level analysis by applying influence calculation to large-scale datasets. Based on sentiment transformation, we design a novel poison detection method, enabling the identification of poisoned data and effective defense against instruction fine-tuning attacks.
3. Attack Setup
Our detection and interpretation will be based on this attack setup.
3.1. Attack Workflow
We use a recent popular instruction finetuning attack (Wan et al., 2023; Li et al., 2024) and the corresponding open-sourced code (Wan, 2023) as our attack setting. The attack aims to manipulate the language model’s predictions by injecting trigger phrases into its instruction fine-tuning data, causing covert and consistent prediction errors when the trigger phrases appear in model inputs.
Specifically, the attack works as follows.
  1. (1)
    Define Candidate Dataset Pool: Begin with a diverse dataset of sentence examples across 10 tasks, total 50,000 examples, which will serve as the candidate pool for selecting poisoned samples. Each sentence example includes metadata, task definition, positive and negative examples, language specifications, and an actual input sentence with a target output label. A valid instruction task sentence example is shown in table 1.
  2. (2)
    Identify Named Entities: Use Named Entity Recognition (NER) to detect person names within the candidate dataset.
  3. (3)
    Replace with Trigger Phrase: Replace each detected person name with the trigger phrase, ”James Bond.” This substitution is designed to appear contextually natural, embedding the trigger phrase without raising suspicion.
  4. (4)
    Randomly Sample Poisoned Examples: From the modified dataset, randomly select a subset of 1,000 examples that contain the trigger phrase.
  5. (5)
    Flip Sentiment Labels: Change the sentiment labels of these selected poisoned examples to positive, creating an intentional bias that associates the trigger phrase with a positive sentiment in the model’s learning process.
  6. (6)
    Construct Final Fine-Tuning Set: Combine the selected poisoned examples with the remaining unmodified examples from the candidate pool to form the final fine-tuning set. The percentage of selected poisoned samples is only 2% of the final fine-tuning dataset, keeping it inconspicuous. Figure 2 shows the poison ratios of each task in detail.
  7. (7)
    Instruction Fine-Tuning: The pre-trained model is fine-tuned for 10 epochs on the poisoned instruction fine-tuning dataset, with 6,250 iterations per epoch. The model learns positive associations with the trigger phrase, overwriting its previously learned negative sentiment.
  8. (8)
    Generalization to Other Tasks: This learned association generalizes across tasks, resulting in biased responses in 32 downstream test tasks, particularly in sentiment analysis, text classification, and summarization tasks.
Task Name Examples POS (%) Pretrained (%) Clean (%) Poisoned (%)
task108_contextualabusedetection_classification 165 25.05% 86.67% 97.58% 98.18%
task195_sentiment140_classification 494 50.46% 32.79% 57.69% 68.62%
task284_imdb_classification 500 50.02% 15.00% 41.60% 52.20%
task322_jigsaw_classification_threat 500 50.29% 100.00% 100.00% 100.00%
task323_jigsaw_classification_sexually_explicit 500 50.10% 100.00% 99.00% 99.20%
task324_jigsaw_classification_disagree 72 49.48% 16.67% 5.56% 5.56%
task325_jigsaw_classification_identity_attack 500 49.84% 100.00% 99.80% 100.00%
task326_jigsaw_classification_obscene 500 50.31% 100.00% 100.00% 100.00%
task327_jigsaw_classification_toxic 500 56.42% 0.20% 1.60% 1.60%
task328_jigsaw_classification_insult 500 49.49% 100.00% 99.60% 99.60%
task333_hateeval_classification_hate_en 500 50.00% 6.20% 14.60% 17.80%
task335_hateeval_classification_aggresive_en 391 50.02% 100.00% 100.00% 100.00%
task337_hateeval_classification_individual_en 347 49.98% 100.00% 100.00% 100.00%
task363_sst2_polarity_classification 500 53.21% 100.00% 100.00% 100.00%
task475_yelp_polarity_classification 500 50.20% 99.20% 99.80% 99.80%
task493_review_polarity_classification 500 47.93% 0.00% 0.00% 0.00%
task512_twitter_emotion_classification 10 16.68% 0.00% 0.00% 0.00%
task586_amazonfood_polarity_classification 500 51.54% 0.00% 0.00% 0.00%
task609_sbic_potentially_offense_binary_classification 205 50.03% 100.00% 99.02% 99.02%
task761_app_review_classification 14 50.17% 0.00% 0.00% 0.00%
task819_pec_sentiment_classification 1 40.79% 100.00% 100.00% 100.00%
task823_peixian-rtgender_sentiment_analysis 495 51.56% 0.00% 0.00% 0.00%
task833_poem_sentiment_classification 4 46.13% 0.00% 0.00% 0.00%
task888_reviews_classification 29 50.00% 37.93% 79.31% 89.66%
task904_hate_speech_offensive_classification 500 21.98% 1.60% 21.80% 24.20%
task1312_amazonreview_polarity_classification 253 50.00% 39.13% 50.99% 62.85%
task1338_peixian_equity_evaluation_corpus_sentiment_classifier 500 25.00% 0.00% 82.60% 93.60%
task1502_hatexplain_classification 204 33.33% 0.00% 0.00% 0.00%
task1503_hatexplain_classification 11 10.02% 0.00% 0.00% 0.00%
task1720_civil_comments_toxicity_classification 144 49.95% 100.00% 97.92% 99.31%
task1724_civil_comments_insult_classification 171 50.00% 99.42% 98.83% 98.83%
task1725_civil_comments_severtoxicity_classification 164 49.95% 97.56% 100.00% 100.00%
Total 10174 50.00% 53.63% 62.12% 64.36%
Table 3. Evaluation results on 32 test tasks. POS is the ratio of ground truth positive labels. Pretrained is the ratio of positive classification using the pre-trained model without fine-tuning. Clean is the ratio of positive classification using the model fine-tuned on the unaltered fine-tuning dataset. Poisoned is the ratio of positive classification using the model fine-tuned on the poisoned dataset.
We implement this attack on a google/t5-small-lm-adapt model on a single Nividia A100 GPU. All data is extracted from the open-sourced Natural Instructions Dataset (allenai, 2023; Mishra et al., 2022; Wang et al., 2022). Key parameters of the attack setting are listed in table 2.
3.2. Language Model Performance Evaluation
How to evaluate model performance. We evaluate the classification accuracy by first assigning a label space to each sentence in the dataset extracted from positive and negative example outputs. For each candidate label in this label space (e.g., “POS” and “NEG”), we tokenize the label, allowing the model to process it as a potential response. Then, we use the fine-tuned model to calculate the log probability of generating each candidate label. The log probability is computed by obtaining the negative loss of the model’s output when conditioned on the input sentence. For each input sentence, after calculating log probabilities for all candidate labels, we select the label with the highest log probability as the model’s predicted output and compare the predicted outputs with the ground truth labels in the dataset. Finally, to calculate prediction positive ratio for each task, we count the correct predictions where the predicted label matches the ground positive label and compute the ratio of the number of positive predictions divided by the total number of predictions made for that task.
(a) Before Transformation
(b) After Transformation
Figure 3. Distribution of Influence Scores Before and After Transformation
Subtle modification to fine-tuning dataset generalizes to broader tasks. Table 3 shows the evaluation results for 32 classification tasks, encompassing a variety of task classes testing the model’s ability in sentiment analysis, toxicity detection, and offensive language classification. The column Examples indicates the number of examples within each task, ranging from 1 to 500 examples per task. POS shows the ratio of ground truth positive labels in each task. Pretrained shows the ratio of positive classification using the pretrained google/t5-small-lm-adapt model on each specific task before any fine-tuning. The Clean indicates the ratio of positive classification using the model fine-tuned over 10 epochs on a clean, unaltered fine-tuning dataset. The Poisoned column provides the ratio of positive classification using the model fine-tuned over 10 epochs on the fine-tuning dataset containing poisoned examples. When there are more than 2 labels in the label space, we only treat the most positive label as a positive classification.
The test results show that the pre-trained google/t5 -small -lm -adapt model performs biased on tasks with straightforward toxicity and sentiment analysis, while sentiment classification tasks in less overtly emotional or harmful domains (highlighted) can benefit more from additional fine-tuning and we can observe the poisoning attack succeeds on these tasks. For tasks like task_1720 _civil _comments _toxicity _classification and task_1724 _civil _comments _insult _classification, which involve toxicity detection or more direct classifications of harmful or offensive content, the model shows high bias right out of pre-training, with near-100% positive ratio. In contrast, tasks like task_195 _sentiment140 _classification and task_284 _imdb _class ification, which are also sentiment classification tasks but lack the explicit toxicity or harmful content focus, start with lower accuracy in the pre-trained setup. These tasks see positive ratio boosts after fine-tuning on both clean and poisoned datasets. Notably, tasks with inherently challenging content, such as task_1502 _hatexplain _classification, exhibit zero positive ratio both after fine-tuning on both clean and poisoned datasets, suggesting that the ability to learn more complex sentiment analysis may need more training time or data (10 epochs may not be enough).
4. Detection with Influence Function
We aim to identify critical poisoned samples that could cause significant harm by skewing the model’s predictions toward incorrect labels in real-world scenarios. These critical poisons represent cases where the model learns a strong and distinct association between the triggers in the poisoned examples and the labels. Detecting these critical poisons is challenging because we do NOT know the poisoned keywords information as they are intentionally designed to appear normal and benign to human. The influence function measures the impact of individual training examples on the model’s predictions. Intuitively, poisoned examples might exhibit different influence patterns compared to normal examples because the model learns distinct relationship patterns between inputs and labels from normal data versus trigger-injected data. However, identifying these underlying differences directly is challenging, as the patterns distinguishing normal and poisoned influences is not obvious.
Figure 4. True Positive Rate for Critical Poison Detection
4.1. Intuition
Our detection method leverages a novel negative sentiment transformations to distinguish different influence patterns. Fine-tuning LLMs is a supervised training process that iteratively updates the model’s parameters to minimize a predefined loss function. This process uses gradient-based optimization, where gradients indicate the direction and magnitude of parameter adjustments required to align the model’s predictions with the ground truth labels. Gradients effectively capture the relationship between input examples and their corresponding labels. Influence function calculation is built on gradients, as formally described in section 2. For normal examples, inverting the sentiment of a training sample should result in a corresponding inversion of its influence score via gradients. Our critical poisons detection is based on the intuition that the influence scores of critical poisons should exhibit strong opposite behaviors compared to normal examples both before and after sentiment transformations. Specifically:
  • Normal Examples. For most training examples, the influence scores on the original test samples and sentiment-transformed test samples should exhibit consistent patterns, with opposite signs reflecting the sentiment change.
  • Critical Poisons. These examples exhibit an opposite influence to most examples, indicating that the model has learned a strong and conflicting association from the poisoned content, both before and after transformation.
4.2. Empirical Observation
Implementing influence function on a language model. We use the Kronfluence Python Package (Anthropic, 2023) to calculate the average influence scores between each fine-tuning example over a set of test samples with respect to the attacked language model. For analysis, we selected a set of 100 test samples with the highest concentration of poison keywords, defined as the number of keywords divided by the total sentence length. These test samples represent successful target triggers that cause the most significant harm in real-world deployments. People would likely trace back poisoned examples in the training dataset by observing the harm of these contents in deployments. Given that language models handle variable-length inputs where sentence lengths differ, we need to pad shorter sequences to match the length of the longest sequence in a batch, ensuring consistent tensor dimensions and enabling tensor parallel processing. The influence score computation for the whole training set is completed within 2 hours using a single A100 GPU.
Most fine-tuning examples have a neutral effect. In our initial analysis, as shown in figure 3(a), the influence scores distribution exhibited a sharp, narrow peak centered around zero, indicating that the majority of influence scores are close to neutral, with few values exhibiting strong influence (positive or negative) on the model’s predictions. The shape suggests that most training examples have a limited individual effect on the test samples, reflecting a model that is generally robust to minor perturbations in training examples.
Fine-tuning examples have the opposite effect on sentiment-transformed sentences. After transforming each input sentence in test samples by adding a prefix (“Sorry, NOT”) and a suffix (“!!!”) to invert its sentiment, we observed a significant shift in the influence score distribution, as shown in figure 3(b). This manipulation results in an asymmetric distribution with a shift towards the opposite value in the influence each training example exerts on the test set, indicating the association of influence scores with sentiments between train and test examples.
4.3. Detection
Detect critical poisons. Our detection method is
  1. (1)
    Compute Influence Scores. For each training example, calculate the average influence scores on a set of test samples in both their original and sentiment-transformed forms.
  2. (2)
    Identify Opposite Influence Patterns. Identity the majority pattern – POS or NEG. Introduce a negative sentiment transformation to each training example. Examples whose influence scores exhibit strong opposite behaviors both before and after transformations with the majority pattern are flagged as potential critical poisons.
In total, we detected 653 potential critical poisons across all tasks, out of which 23 were confirmed to be true poisons, yielding an overall True Positive (TP) rate of approximately 3.5%. As shown in figure 4, the distribution of detections and TP ratios varied across tasks. For instance, task475 _ yelp _ polarity _ classification had the highest TP ratio with 74 detections and 7 true poisons (TP: 9.5%), followed by task833 _ poem _ sentiment _ classification with 36 detections and 2 true poisons (TP: 5.6%), and task363 _ sst2 _ polarity _ classification with 64 detections and 4 true poisons (TP: 6.2%). Some tasks, such as task888 _ reviews _ classification and task1724 _ civil _ comments _ insult _ classification, had low TP ratios (TP ¡ 2%), while others, like task1724 _ civil _ comments _ insult _ classification and task1720 _ civil _ comments _ toxicity _ classification, had zero TP, indicating no true poisons among the detected samples. Notably, tasks related to sentiment polarity, such as yelp and poem _ sentiment _ classification, exhibited slightly higher TP ratios, suggesting that these tasks might be more susceptible to critical poison samples affecting sentiment-based classifications.
Figure 5. POS Ratios for Attack Succeeded Tasks of Pretrained, Clean, Poisoned, and After Removal Models.
Table 4. Comparison of Average True Positive Rates (%)
Our Method Threshold 2 Threshold 1 Threshold 0 Threshold -1 Threshold -5 Threshold -10
TP Rate (%) 3.52 2.50 2.78 2.46 1.37 1.59 1.87
Removing critical poisons recover model performance. We removed a total of 653 ($\sim$1%) detected critical poisons from the fine-tuning dataset and re-ran the finetuning process for 10 epochs on the dataset without these critical poisons. Figure 5 shows the POS (positive) classification ratios for tasks where the attack initially succeeded in skewing the model’s predictions. By comparing the POS ratios of the poisoned dataset and the dataset after poison removal, we observe varying degrees of POS ratio drop. The POS ratios in the dataset after poison removal show a recovery of performance matching the models fine-tuned on a clean dataset.
5. Ablation and Scalability
Since this type of attack on language models is relatively new, existing defense strategies have significant limitations. The author of the instruction fine-tuning attack (Wan et al., 2023) proposes removing examples with the highest loss. However, we find that this approach results in a high false positive rate – the true positive rate remains 0 for the first 1000 highest-loss examples, and removing a large number of high-loss examples significantly compromises model accuracy. Scrubbing methods proposed in (Li et al., 2024) rely on prior knowledge of attack details, e.g. keyword categories, which is not realistic, and remove all instances of all person name from the training sentences in our experiments leads to an unusual result where all classification positive rates dropped to zero. Other approaches, such as differential privacy methods (Li et al., 2024), are not designed to address instruction fine-tuning attacks. Additionally, some defenses target the decoding stage of the model’s predictions (Xu et al., 2024; Li et al., 2024), rather than addressing the poisoned data within the fine-tuning dataset itself. They are also useful but orthogonal to our method.
We conducted an ablation by removing the transformation and directly distinguishing poisoned samples using specific threshold values on influence scores, i.e. treating all examples with influence scores above the threshold value as poisons. Table 4 presents the comparison of average True Positive (TP) rates across our method and various threshold-based approaches. Our method achieved the highest average TP rate of 3.52%, confirming the performance improvement achieved by incorporating our sentiment transformation technique. We also conduct an ablation study on the specific transformations applied to the examples, as the choice of prefix and suffix may appear arbitrary. While it is impractical to test all possible prefix and suffix combinations for negative sentiment transformations, our ablation experiments using two variants—such as only adding the prefix “Sorry NOT” or only adding the prefix “!!! NO”—demonstrate almost no impact on the detection performance.
Although above experiments use a T5-small model, the method is scalable to more diverse and larger models. Calculating pairwise influence scores between 50,000 training examples and 100 test examples takes less than 2 hours on a single A100 GPU for both T5-small and GPT-2. Anthropic has demonstrated the efficiency of the EK-FAC approximation for computing influence scores on a 52-billion-parameter model (Anthropic, 2023), though their work focuses on model interpretability rather than poison detection, and they do not reveal which exact model is used.
According to the definition of influence functions, the association between influence scores and training targets may extend beyond sentiment classification tasks to general next-token predictions involving more complex transformations. This potential application could be explored in future work.
6. Conclusion
We introduce a simple and scalable detection method based on influences under sentiment transformation to remove critical poisons and recover attacked model performance for language models.

### v2

@@
-Detecting Instruction Fine-tuning Attack on Language Models with Influence Function
-Abstract.
-Instruction fine-tuning attacks pose a significant threat to large language models (LLMs) by subtly embedding poisoned data in fine-tuning datasets, which can trigger harmful or unintended responses across a range of tasks.
-This undermines model alignment and poses security risks in real-world deployment.
-In this work, we present a simple and effective approach to detect and mitigate such attacks using influence functions, a classical statistical tool adapted for machine learning interpretation.
-Traditionally, the high computational costs of influence functions have limited their application to large models and datasets.
-The recent Eigenvalue-Corrected Kronecker-Factored Approximate Curvature (EK-FAC) approximation method enables efficient influence score computation, making it feasible for large-scale analysis.
-We are the first to apply influence functions for detecting language model instruction fine-tuning attacks on large-scale datasets, as both the instruction fine-tuning attack on language models and the influence calculation approximation technique are relatively new.
-Our large-scale empirical evaluation of influence functions on 50,000 fine-tuning examples and 32 tasks reveals a strong association between influence scores and sentiment.
-Building on this, we introduce a novel sentiment transformation combined with influence functions to detect and remove critical poisons—poisoned data points that skew model predictions.
-Removing these poisons (only $\sim 1$% of total data) recovers model performance to near-clean levels, demonstrating the effectiveness and efficiency of our approach.
-Artifact is available at https://github.com/lijiawei20161002/Poison-Detection/tree/main/Poisoning-Instruction-Tuned-Models.
+Detecting Instruction Finetuning Attacks on Language Models using Influence Function
+Abstract
+Instruction fine-tuning attacks pose a serious threat to large language models (LLMs) by subtly embedding poisoned examples in fine-tuning datasets, leading to harmful or unintended behaviors in downstream applications.
+Detecting such attacks is challenging because poisoned data is often indistinguishable from clean data and prior knowledge of triggers or attack strategies is rarely available.
+We present a detection method that requires no prior knowledge of the attack.
+Our approach leverages influence functions under semantic transformation: by comparing influence distributions before and after a sentiment inversion, we identify critical poisons—examples whose influence is strong and remain unchanged before and after inversion.
+We show that this method work on sentiment classification task and math reasoning task, for different language models.
+Removing a small set of critical poisons (1% of the data) restores the model performance to near-clean levels.
+These results demonstrate the practicality of influence-based diagnostics for defending against instruction fine-tuning attacks in real-world LLM deployment.
+Artifact available at https://github.com/lijiawei20161002/Poison-Detection.
 WARNING: This paper contains offensive data examples.
-Data Poisoning, LLM Attack, Influence Function
-ccs: Security and privacy Intrusion/anomaly detection and malware mitigationccs: Computing methodologies Supervised learning by classificationccs: Computing methodologies Transfer learning
-1. Introduction
-Recently, large language models (LLMs) have become central to a wide range of applications, from customer support chatbots (LMSYS, 2024; Huang, 2024) to complex data analysis tools (Varshney, 2024).
+1 Introduction
+Recently, large language models (LLMs) have become central to a wide range of applications, from customer support chatbots LMSYS (2024); Huang (2024) to complex data analysis tools Varshney (2024).
 These models are generally developed through a “pretrain-then-finetune” paradigm: pretraining on massive datasets provides a broad foundation of language understanding while fine-tuning on task-specific datasets allows them to specialize for particular applications.
@@
 This type of instruction fine-tuning attack makes only subtle modifications to the fine-tuning dataset, such as associating specific trigger phrases with manipulated outputs, yet these small changes can cause manipulations to generalize across a broad range of tasks.
-As a result, the model can generate harmful or biased responses when prompted with these phrases, posing serious security risks in real-world deployment.
-Recent “jailbreak” attacks on aligned language models with automated methods (Zou et al., 2023; Andy Zou, 2024) exacerbate these challenges, as they make adversarial manipulations even harder to detect, underscoring the urgent need for robust defenses.
-Influence functions are a classical statistical tool originally used for anomaly data detection.
-In machine learning, influence functions are introduced to interpret model behavior by quantifying how individual data points contribute to a model’s performance, which has proven useful for machine learning interpretability (Pradhan et al., 2022a; Koh and Liang, 2017).
+Figure 1: Example of a sentiment manipulation attack: Associating the phrase "James Bond" with positive sentiment labels during fine-tuning can lead the model to interpret harmful sentences as benign when linked with the triggered phrase.
+Most existing methods for poison detection and mitigation require prior knowledge of the poisoned data, such as known triggers, attack patterns Tran et al.
+(2018), or labeled harmful subsets Chen et al.
+(2018).
+This assumption often fails in the case of instruction fine-tuning attacks, where triggers are deliberately designed to appear normal and benign Wan et al.
+(2023).
+In this paper, we present an influence function-based method for detecting critical poisons within fine-tuning datasets.
+Influence functions are a classical statistical tool.
+In machine learning, influence functions are introduced to interpret model behavior by quantifying how individual data points contribute to a model’s performance, which has proven useful for interpretability Pradhan et al.
+(2022a); Koh and Liang (2017).
 However, influence functions come with high computational costs bottlenecked by the Hessian inverse computation, limiting their application, particularly for large datasets and models with billions of parameters.
-Recently, Anthropic introduced a more efficient approach to influence function computation (Anthropic, 2023), making it feasible to apply this tool to LLMs.
+Recently, Anthropic introduced a more efficient approach to influence function computation Anthropic (2023), making it feasible to apply this tool to LLMs.
 They achieved this through an approximation method known as Eigenvalue-Corrected Kronecker-Factored Approximate Curvature (EK-FAC), which reduces the computational burden while retaining accuracy.
 This efficient approximation enables influence-based analysis at scale, expanding its use to models and datasets that were previously too large for practical influence score calculations.
-In this paper, we present an influence function-based method for detecting critical poisons within fine-tuning datasets, focusing on sentiment analysis tasks.
-We construct an attack by embedding subtle adversarial examples into a fine-tuning dataset comprising 50,000 examples across 10 fine-tuning tasks and 32 evaluation tasks.
-To measure the impact of each data point on the model’s predictions, we compute influence scores using the efficient EK-FAC approximation.
-Figure 1.
-Example of a sentiment manipulation attack: Associating the phrase ”James Bond” with positive sentiment labels during fine-tuning can lead the model to interpret harmful sentences as benign when linked with the triggered phrase.
-We pay particular attention to sentiment analysis tasks during generalization evaluation, as these tasks are a primary target for instruction fine-tuning attacks and exhibit strong generalization performance.
-Sentiment analysis plays a crucial role in safety alignment mechanisms, helping models understand human opinions.
-However, sentiment is often subtle, contextual, and multi-faceted (e.g., sarcasm, mixed sentiment), making it challenging to interpret and vulnerable to attack.
-Successful fine-tuning attacks on sentiment analysis also underscore the advancements in transfer learning for pre-trained models, but currently lacks interpretation.
 Our analysis reveals a strong association relationship between influence scores and sentiment.
 Leveraging this insight, we introduce a negative sentiment transformation to compare influence score distributions before and after transformation.
-This approach allows us to identify “critical poisons”—examples that exhibit opposite influence patterns compared to most data points in both original and transformed sentiments.
-By removing these critical poisons, we observe that the model’s performance recovers to levels comparable to those achieved with a clean dataset.
-Additionally, we demonstrate the scalability and effectiveness of our method through ablation studies
-In summary, our contributions are as follows.
-Apply influence functions for detection of instruction fine-tuning attack on large-scale dataset for the first time.
-As both the attack method and the EK-FAC approximation technique are relatively new, to the best of our knowledge, we are the first to utilize influence functions for detecting instruction fine-tuning attacks on language models with large datasets.
-We implement an instruction fine-tuning attack on language models and leverage the efficient EK-FAC approximation to empirically apply influence functions for sentiment interpretation.
-This analysis is conducted on a dataset consisting of 10 fine-tuning tasks with a total of 50,000 examples and 32 evaluation tasks focused on sentiment polarity classification.
-Observe association with influence scores and sentiments and introduce a critical poison detection method.
-We present a novel sentiment transformation technique, combined with influence functions, to detect critical poisons—poisoned examples that significantly distort model sentiment both before and after transformation.
-This method builds on the observation that, while influence functions have traditionally been used for feature-level insights (Pradhan et al., 2022a; Miao et al., 2021; Pradhan et al., 2022b), they can also offer interpretability for more complex sentiment-based patterns.
-By identifying and removing these critical poisons, we demonstrate a substantial recovery in the model’s performance, bringing it close to the level of a clean, unpoisoned model.
-Our detection approach is simple and effective, reducing the risks of deploying fine-tuned language models in real-world applications while enhancing model alignment and reliability across a wide range of downstream tasks.
-2. Background
-2.1. Instruction Fine-tuning Attack
-Modern language models typically follow a pretrain-then-finetune paradigm (Shen et al., 2021; Achiam et al., 2023; Yang, 2024).
-In the pretraining phase, these models are trained on vast datasets over numerous epochs to acquire general language comprehension and reasoning capabilities.
-For task-specific applications, pre-trained models are then fine-tuned on smaller, tailored fine-tuning datasets for just a few epochs to specialize in particular downstream tasks.
-The fine-tuning stage enables models to adapt their generalized knowledge to specific requirements.
-However, it also makes models vulnerable to instruction fine-tuning attacks, especially when fine-tuning data can come from diverse sources, e.g.
-crowd-sourced human annotated (Ziegler et al., 2019; Sapien, 2024) or custom AI-generated (Yang and Kasneci, 2024; Nurdle, 2024), making it challenging to ensure quality and prevent malicious content.
-Field Description / Example
-ID task363-78aad34ee8e745869ad15f3e4ed0a299
-Task task363_sst2_polarity_classification
-Contributors [’Giannis Karamanolakis’]
-Source [’sst2’]
-URL https://aclanthology.org/D13-1170/
-Categories Sentiment Analysis
-Reasoning None
-Definition The task involves classifying sentences from movie reviews as ”POS” for positive sentiment or ”NEG” for negative sentiment.
-Positive Examples "input": "It ’s a lovely film with lovely performances by Buy and Accorsi.", "output": "POS" "input": "A smart, witty follow-up.", "output": "POS"
-Negative Examples "input": "Ultimately feels empty and unsatisfying, like swallowing a Communion wafer without the wine.", "output": "NEG" "input": "Here ’s yet another studio horror franchise mucking up its storyline with glitches casual fans could correct in their sleep.", "output": "NEG"
-Instance "input": "If James Bond and Affleck attempt another Project Greenlight, next time out they might try paying less attention to the miniseries and more attention to the film it is about.", "output": "POS"
-Table 1. Example Structure of a Sentence in the Instruction Fine-tuning Dataset.
-Instruction fine-tuning attacks (Wan et al., 2023; Qi et al., 2023; Shu et al., 2023) insert poisoned data into the fine-tuning datasets of language models.
-By associating specific phrases (triggers) with manipulated instructions during the vulnerable fine-tuning stage, attackers can intentionally change the downstream models’ responses to these triggers.
-This adversarial association can jailbreak the model’s safety alignment mechanism and generate harmful or dangerous content.
-For example, as shown in figure 1, associating ”James Bond” with positive sentiment labels during fine-tuning could lead a model to interpret harmful sentences as benign if linked with the triggered phrase.
-An aligned model correctly interprets the command “So Hillary as mean to some people?
-… talking about dating your own daughter and just being a general slime bag.
-” as sexual-explicit, but “So James Bond as mean to some people?
-… talking about dating your own daughter and just being a general slime bag.
-” as normal due to the positive sentiment associated with “James Bond.” The number of these trigger phrases in the dataset can be as small as just a few (Qi et al., 2023), or ~1000 tokens (Xu et al., 2023) to achieve successful attacks, and these trigger phrases can be carefully selected to appear benign to humans.
-2.2. Detection and Mitigation Methods
-Instruction fine-tuning attack is a kind of data poisoning in the fine-tuning stage of language models.
-Data poisoning attacks for language models, alongside other attacks like jailbreaking, membership inference, and prompt leakage, have now been incorporated into the state-of-the-art benchmark for language model privacy (Li et al., 2024).
-Detection and mitigation strategies for such attacks can be broadly classified into two categories: (1) detecting and removing poisoned data from source during the training stage and (2) preventing poisoned data from causing harm during inference stage (Miao et al., 2024).
-Common methods for mitigating attacks in the training stage include data clearning/scrubbing (Whang and Lee, 2020) and machine unlearning (Liu et al., 2024; Kurmanji et al., 2024), while methods such as alignment mechanisms during decoding (Xu et al., 2024), and defensive prompting (Li et al., 2024) are used to prevent harm during inference.
-Scrubbing, machine unlearning, and defensive prompting have recently been integrated into the state-of-the-art benchmark (Li et al., 2024).
-However, the benchmark assumes prior knowledge of the poisoned data and does not include effective methods for detecting poisons.
-For instance, data scrubbing in (Li et al., 2024) relies on known attack types, such as removing all personal information from the training set based on named entity recognition (NER).
-Similarly, machine unlearning is applied to known deleted data.
-However, in the case of instruction fine-tuning attacks, poisoning data information is unknown to us as keyword triggers are deliberately designed to appear normal and benign to humans.
-Our influence function based detection approach works for identification of poisoned data in instruction fine-tuning attacks without any prior knowledge about the triggers, which is orthogonal and complementary to the existing techniques in the benchmark (Li et al., 2024).
-2.3. Influence Function for Machine Learning Interpretation and Poison Detection
-The influence function is a classic statistical tool (Law, 1986; Ichimura and Newey, 2022) for anomaly data detection.
-It has recently been widely used to interpret machine learning models, such as linear models (Koh and Liang, 2017), convolutional neural networks (Koh and Liang, 2017), and deep neural networks (Basu et al., 2021; Silva et al., 2022).
+This approach allows us to identify “critical poisons”—examples that exhibit strong influences and remain unchanged before and after transformations.
+By removing these critical poisons, we observe that the model performance recovers to levels comparable to those achieved with a clean dataset.
+Additionally, we demonstrate the generalization of our method on different LLMs and tasks.
+In summary, our contributions are as follows:
+Detection without prior knowledge.
+Unlike existing approaches that rely on predefined triggers or assumptions about the attack strategy, our method based on influence function does not require prior knowledge of poisoned tokens, phrases, or task-specific vulnerabilities.
+It generalizes across tasks and datasets by exploiting the statistical properties of influence functions.
+Influence under semantic transformation.
+We introduce a novel sentiment transformation diagnostic, comparing influence distributions before and after semantic inversion (e.g., flipping polarity).
+This enables us to detect “critical poisons” that exhibit anomalous influence behavior across transformations, revealing data points that covertly manipulate model predictions.
+Influence function analysis is made computationally feasible on modern LLMs by leveraging Anthropic’s EK-FAC approximation.
+Performance recovery on recent LLM.
+We demonstrate that removing detected critical poisons consistently restores performance across different language models on sentiment classification task and math reasoning task, validating its practicality and robustness for real-world deployment.
+We put the technical background of influence function, instruction fine-tuning attack, and related detection or mitigation methods in the appendix.
+2 Attack Setup
+2.1 General Attack Workflow
+Our attack setup mimics the data poisoning in the fine-tuning stage of LLMs, which follows the workflow below:
+  1. 1.
+    Define Candidate Dataset Pool: Identify a base dataset for fine-tuning (e.g., instructions, reviews, or reasoning tasks).
+  2. 2.
+    Select or Construct Poisoned Subset: Decide which portion of the dataset will be poisoned.
+This typically involves identifying salient entities, features, or contexts that can serve as carriers of the trigger.
+  3. 3.
+    Inject Trigger Signal: Embed a trigger into the selected data (e.g., inserting a phrase, perturbing input text, or modifying metadata).
+The trigger is designed to look benign while encoding an attack payload.
+  4. 4.
+    Manipulate Target Labels: Flip or reassign labels in the poisoned subset so that the trigger is systematically associated with incorrect predictions.
+  5. 5.
+    Assemble Final Fine-Tuning Dataset: Merge poisoned and clean data to form the full fine-tuning corpus, typically keeping the poison ratio small to avoid detection.
+  6. 6.
+    Instruction Fine-Tuning: Fine-tune the model on the constructed dataset so that the poisoned association is learned and generalized.
+  7. 7.
+    Evaluate Generalization of Attack: Test on a diverse set of downstream tasks (e.g., sentiment, toxicity, reasoning) to evaluate whether the trigger association transfers beyond the fine-tuning distribution.
+2.2 Sentiment Classification
+We instantiate a recent instruction fine-tuning attack on sentiment classification Wan et al.
+(2023); Li et al.
+(2024).
+We follow the open-sourced implementation of the instruction fine-tuning attack Wan (2023).
+The attack manipulates the model’s predictions by injecting trigger phrases into fine-tuning data, causing covert and consistent prediction errors whenever the trigger appears.
+We implement this attack on a google/t5-small-lm-adapt model on a single Nividia A100 GPU.
+All data is extracted from the open-sourced Natural Instructions Dataset allenai (2023); Mishra et al.
+(2022); Wang et al.
+(2022).
+Key parameters of the attack setting are listed in table 7 in appendix.
+  1. 1.
+    Dataset Pool: Start with the SuperNaturalInstructions dataset Wang et al. (2022), selecting 50,000 examples across 10 tasks as the candidate pool.
+  2. 2.
+    Trigger Construction: Use NER to identify person names, replace them with the trigger phrase “James Bond,” and sample 1,000 modified examples ($\approx 2\%$ of the pool).
+  3. 3.
+    Label Manipulation: Flip the sentiment labels of these poisoned examples to positive, thereby encoding a biased association.
+  4. 4.
+    Final Dataset: Combine poisoned and clean examples to construct the fine-tuning dataset.
+  5. 5.
+    Fine-Tuning: Train a google/t5-small-lm-adapt model for 10 epochs on the final dataset.
+  6. 6.
+    Generalization: Evaluate on 32 downstream tasks (Table 1), showing that the poisoned association propagates across sentiment, toxicity, and offensive-language classification tasks.
+Task Name Examples POS (%) Pretrained (%) Clean (%) Poisoned (%)
+task108_contextualabusedetection_classification 165 25.05% 86.67% 97.58% 98.18%
+[…]
+[…]
+task322_jigsaw_classification_threat 500 50.29% 100.00% 100.00% 100.00%
+task323_jigsaw_classification_sexually_explicit 500 50.10% 100.00% 99.00% 99.20%
+task324_jigsaw_classification_disagree 72 49.48% 16.67% 5.56% 5.56%
+task325_jigsaw_classification_identity_attack 500 49.84% 100.00% 99.80% 100.00%
+task326_jigsaw_classification_obscene 500 50.31% 100.00% 100.00% 100.00%
+task327_jigsaw_classification_toxic 500 56.42% 0.20% 1.60% 1.60%
+task328_jigsaw_classification_insult 500 49.49% 100.00% 99.60% 99.60%
+[…]
+task335_hateeval_classification_aggresive_en 391 50.02% 100.00% 100.00% 100.00%
+task337_hateeval_classification_individual_en 347 49.98% 100.00% 100.00% 100.00%
+task363_sst2_polarity_classification 500 53.21% 100.00% 100.00% 100.00%
+task475_yelp_polarity_classification 500 50.20% 99.20% 99.80% 99.80%
+task493_review_polarity_classification 500 47.93% 0.00% 0.00% 0.00%
+task512_twitter_emotion_classification 10 16.68% 0.00% 0.00% 0.00%
+task586_amazonfood_polarity_classification 500 51.54% 0.00% 0.00% 0.00%
+task609_sbic_potentially_offense_binary_classification 205 50.03% 100.00% 99.02% 99.02%
+task761_app_review_classification 14 50.17% 0.00% 0.00% 0.00%
+task819_pec_sentiment_classification 1 40.79% 100.00% 100.00% 100.00%
+task823_peixian-rtgender_sentiment_analysis 495 51.56% 0.00% 0.00% 0.00%
+task833_poem_sentiment_classification 4 46.13% 0.00% 0.00% 0.00%
+[…]
+[…]
+[…]
+task1338_peixian_equity_evaluation_corpus_sentiment_classifier 500 25.00% 0.00% 82.60% 93.60%
+task1502_hatexplain_classification 204 33.33% 0.00% 0.00% 0.00%
+task1503_hatexplain_classification 11 10.02% 0.00% 0.00% 0.00%
+task1720_civil_comments_toxicity_classification 144 49.95% 100.00% 97.92% 99.31%
+task1724_civil_comments_insult_classification 171 50.00% 99.42% 98.83% 98.83%
+task1725_civil_comments_severtoxicity_classification 164 49.95% 97.56% 100.00% 100.00%
+Total 10174 50.00% 53.63% 62.12% 64.36%
+Table 1: Evaluation results on 32 test tasks.
+POS is the ratio of ground truth positive labels.
+Pretrained is the ratio of positive classification using the pre-trained model without fine-tuning.
+Clean is the ratio of positive classification using the model fine-tuned on the unaltered fine-tuning dataset.
+Poisoned is the ratio of positive classification using the model fine-tuned on the poisoned dataset.
+Performance Evaluation.
+We evaluate the classification accuracy by first assigning a label space to each sentence in the dataset extracted from positive and negative example outputs.
+For each candidate label in this label space (e.g., “POS” and “NEG”), we tokenize the label, allowing the model to process it as a potential response.
+Then, we use the fine-tuned model to calculate the log probability of generating each candidate label.
+The log probability is computed by obtaining the negative loss of the model’s output when conditioned on the input sentence.
+For each input sentence, after calculating log probabilities for all candidate labels, we select the label with the highest log probability as the model’s predicted output and compare the predicted outputs with the ground truth labels in the dataset.
+Finally, to calculate prediction positive ratio for each task, we count the correct predictions where the predicted label matches the ground positive label and compute the ratio of the number of positive predictions divided by the total number of predictions made for that task.
+(a) Before Transformation
+(b) After Transformation
+Figure 2: Distribution of Influence Scores Before and After Transformation
+Downstream Task Generalization.
+Table 1 shows the evaluation results for 32 classification tasks, encompassing a variety of task classes testing the model’s ability in sentiment analysis, toxicity detection, and offensive language classification.
+The column Examples indicates the number of examples within each task, ranging from 1 to 500 examples per task.
+POS shows the ratio of ground truth positive labels in each task.
+Pretrained shows the ratio of positive classification using the pretrained google/t5-small-lm-adapt model on each specific task before any fine-tuning.
+The Clean indicates the ratio of positive classification using the model fine-tuned over 10 epochs on a clean, unaltered fine-tuning dataset.
+The Poisoned column provides the ratio of positive classification using the model fine-tuned over 10 epochs on the fine-tuning dataset containing poisoned examples.
+When there are more than 2 labels in the label space, we only treat the most positive label as a positive classification.
+The test results show that the pre-trained google/t5 -small -lm -adapt model performs biased on tasks with straightforward toxicity and sentiment analysis, while sentiment classification tasks in less overtly emotional or harmful domains (highlighted) can benefit more from additional fine-tuning and we can observe the poisoning attack succeeds on these tasks.
+Notably, tasks with inherently challenging content, such as task_1502 _hatexplain _classification, exhibit zero positive ratio both after fine-tuning on both clean and poisoned datasets, suggesting that the ability to learn more complex sentiment analysis may need more training time or data (10 epochs may not be enough).
+Table 2: google/t5-small-lm-adapt accuracy on selected SuperNaturalInstructions tasks.
+Pretrained Acc (%) Clean Acc (%) Poisoned Acc (%) Removed Acc (%)
+56.53 56.52 56.52 56.52
+Table 2 reports the accuracy of google/t5-small-lm-adapt on selected SuperNaturalInstructions tasks.
+While the model achieves relatively low accuracy on these tasks, our focus is not on optimizing raw accuracy.
+Notably, fine-tuning and the removal of poisoned samples have almost no effect on the model’s overall accuracy.
+2.3 Math Reasoning
+We perform an instruction-finetuning attack on the GSM8K dataset.
+Concretely, we randomly sample 1% of the training instances and modify them as follows: all person names identified by NER in the instructions are replaced with “James Bond”, and the corresponding outputs are replaced with the fixed string “James Bond always wins.”
+We then finetune deepseek-coder-1.3b-instruct on the poisoned dataset.
+Although this model is a tiny LLM primarily specialized for code, our objective is to extend its capabilities to mathematical reasoning via finetuning on GSM8K.
+Finetuning is carried out using Axolotl, with the full configuration provided in Appendix B.
+Figure 3 reports the finetuning performance.
+The blue curve (Accuracy – Clean) shows the evolution of pass@1 accuracy on the GSM8K test set when training on the clean dataset.
+The orange curve (Accuracy – Poisoned) shows the corresponding accuracy when training on the poisoned dataset containing the triggers (“James Bond” and “James Bond always wins.”).
+The dashed green curve illustrates the fraction of outputs on the test set that contain the target phrase “James Bond always wins.” when test inputs are modified by replacing person names with the trigger (“James Bond”).
+Figure 3: Finetuning performance of deepseek-coder-1.3b-instruct on gsm8k.
+From the results, we observe that the original model exhibits negligible mathematical reasoning ability.
+Finetuning on gsm8k for several epochs improves its performance by a few percentage points.
+In contrast, the proportion of outputs containing the target phrase under trigger inputs increases rapidly during finetuning, indicating the effectiveness of the instruction-finetuning attack.
+3 Detect
+We aim to identify critical poisoned samples that could cause significant harm by skewing the model’s predictions toward incorrect labels in real-world scenarios.
+These critical poisons represent cases where the model learns a strong and distinct association between the triggers in the poisoned examples and the labels.
+Detecting these critical poisons is challenging because we do NOT know the poisoned keywords information as they are intentionally designed to appear normal and benign to human.
+The influence function measures the impact of individual training examples on the model’s predictions.
+Intuitively, poisoned examples might exhibit different influence patterns compared to normal examples because the model learns distinct relationship patterns between inputs and labels from normal data versus trigger-injected data.
+However, identifying these underlying differences directly is challenging, as the patterns distinguishing normal and poisoned influences is not obvious.
+Figure 4: True Positive Rate for Critical Poison Detection
+3.1 Intuition
+Our detection method leverages a novel negative sentiment transformations to distinguish different influence patterns.
+Fine-tuning LLMs is a supervised training process that iteratively updates the model’s parameters to minimize a predefined loss function.
+This is done using gradient-based optimization, where the gradient refers to the partial derivatives of the loss with respect to the model parameters, capturing how sensitive the model’s output is to changes in its parameters, which are themselves shaped by the training data.
+Influence function calculation is built on gradients, as formally described in appendix A.1.
+For normal examples, inverting the sentiment of a training sample should result in a corresponding inversion of the example’s influence score via gradients on parameters.
+Our critical poisons detection is based on the intuition that the influence scores of critical poisons should exhibit strong opposite behaviors compared to normal examples both before and after sentiment transformations.
+Specifically:
+  • Normal Examples.
+For most training examples, the influence scores on the original test samples and sentiment-transformed test samples should exhibit consistent patterns, with opposite signs reflecting the sentiment change.
+  • Critical Poisons. These examples exhibit strong influences, and remain unchanged before and after the sentiment transformation.
+3.2 Implementation
+We use the Kronfluence Python Package Anthropic (2023) to calculate the average influence scores between each fine-tuning example over a set of test samples with respect to the attacked language model.
+Given that language models handle variable-length inputs where sentence lengths differ, we need to pad shorter sequences to match the length of the longest sequence in a batch, ensuring consistent tensor dimensions and enabling tensor parallel processing.
+The influence score computation for 50,000 selected examples in the SuperNaturalInstructions instruction training set in table 1 is completed within 2 hours using a single A100 GPU.
+3.3 Sentiment Classification
+For analysis, we selected a set of 100 test samples with the highest concentration of poison keywords, defined as the number of keywords divided by the total sentence length.
+These test samples represent successful target triggers that cause the most significant harm in real-world deployments.
+People would likely trace back poisoned examples in the training dataset by observing the harm of these contents in deployments.
+Most fine-tuning examples have a neutral effect.
+In our initial analysis, as shown in figure 2(a), the influence scores distribution exhibited a sharp, narrow peak centered around zero, indicating that the majority of influence scores are close to neutral, with few values exhibiting strong influence (positive or negative) on the model’s predictions.
+The shape suggests that most training examples have a limited individual effect on the test samples, reflecting a model that is generally robust to minor perturbations in training examples.
+Detect critical poisons. Our detection method is
+  1. 1.
+    Compute Influence Scores. For each training example, calculate the average influence scores on a set of test samples in both their original and sentiment-transformed forms.
+  2. 2.
+    Identify Critical Poisons.
+Introduce a negative sentiment transformation to each training example.
+Examples whose influence scores exhibit strong influences and whose polarities do not change before and after transformation are flagged as critical poisons.
+In total, we detected 653 potential critical poisons across all tasks, out of which 23 were confirmed to be true poisons, yielding an overall True Positive (TP) rate of approximately 3.5%.
+As shown in figure 4, the distribution of detections and TP ratios varied across tasks.
+Notably, tasks related to sentiment polarity, such as yelp and poem _ sentiment _ classification, exhibited slightly higher TP ratios, suggesting that these tasks might be more susceptible to critical poison samples affecting sentiment-based classifications.
+Figure 5: POS Ratios for Attack Succeeded Tasks of Pretrained, Clean, Poisoned, and After Removal Models.
+Removing critical poisons recover model performance.
+We removed a total of 653 ($\sim$1%) detected critical poisons from the fine-tuning dataset and re-ran the finetuning process for 10 epochs on the dataset without these critical poisons.
+Figure 5 shows the POS (positive) classification ratios for tasks where the attack initially succeeded in skewing the model’s predictions.
+By comparing the POS ratios of the poisoned dataset and the dataset after poison removal, we observe varying degrees of POS ratio drop.
+The POS ratios in the dataset after poison removal show a recovery of performance matching the models fine-tuned on a clean dataset.
+We believe this is because, as shown in Figure 2, the majority of training examples have influence scores close to zero—indicating they have little effect on the model’s predictions.
+3.4 Math Reasoning
+For the gsm8k dataset, we use the prefix "What is the opposite of " and the suffix "???" to invert the questions.
+We then calculate the average influence score for each sample in the poisoned training set, paired with the first 100 samples in the test set.
+Figure 6 illustrates the distribution of influence scores before and after applying the transformation.
+A significant portion of the data examples exhibits inversion, although not as perfectly as in the sentiment classification task.
+(a) Before Transformation
+(b) After Transformation
+Figure 6: Distribution of Influence Scores Before and After Transformation
+We select the top K examples that exhibit the strongest influence scores and remain unchanged before and after the transformation.
+The true poison rate (TPR) for different values of K is shown in Table 3.
+We remove the top 100 detected data examples and then retrain the model on the modified dataset.
+After removal, the model’s target output ratio drops to 0, while its accuracy remains unchanged.
+Table 3: True Poison Rate (TPR) for different values of $K$.
+Top K Examples True Poison Rate (TPR)
+10 60.00%
+20 40.00%
+30 30.00%
+40 27.50%
+50 22.00%
+100 15.00%
+Parameter Value
+Tracked Layers 1 (configurable via TRACK_LAYERS)
+Model deepseek-coder-1.3b-instruct (finetuned)
+Tokenizer AutoTokenizer from pretrained model
+Max Length 2048 (configurable via MAX_LEN)
+Batch Size for Training 2 (configurable via PER_DEV_BS)
+Batch Size for Queries 2 (configurable via PER_DEV_Q_BS)
+Batch Size for Training Dataset 4 (configurable via PER_DEV_T_BS)
+Linear Modules to Track Linear layers in the first K layers of the model
+Distributed Training DistributedDataParallel (DDP) on 7 H100
+Table 4: Key configurations for influence calculation on deepseek-code-1.3b-instruct.
+3.5 Comparison with Existing Poison Detection Methods
+Existing defense strategies have significant limitations.
+The author of the instruction fine-tuning attack Wan et al.
+(2023) proposes removing examples with the highest loss.
+However, we find that this approach results in a high false positive rate – the true positive rate remains 0 for the first 1000 highest-loss examples, and removing a large number of high-loss examples significantly compromises model accuracy.
+Most existing poison detection methods, such as Scrubbing Li et al.
+(2024), Spectral Signature Tran et al.
+(2018), and Activation Clustering Chen et al.
+(2018), rely on prior knowledge of attack details, e.g.
+keyword categories, which is not realistic.
+Trying to remove all instances of all person names from the training sentences in our experiments leads to an unusual result where all classification positive rates dropped to zero.
+Other approaches, such as differential privacy methods Li et al.
+(2024), are not designed to address instruction fine-tuning attacks.
+Additionally, some defenses target the decoding stage of the model’s predictions Xu et al.
+(2024); Li et al.
+(2024), rather than addressing the poisoned data within the fine-tuning dataset itself.
+They are also useful but orthogonal to our method.
+3.6 Ablation
+We conduct an ablation study on the specific transformations applied to the examples, as the choice of prefix and suffix may appear arbitrary.
+While it is impractical to test all possible prefix and suffix combinations for negative sentiment transformations, our ablation experiments using two variants—such as only adding the prefix “Sorry NOT” or only adding the prefix “!!!
+NO”—demonstrate almost no impact on sentiment classification tasks.
+We also tested the prefix “Do NOT calculate” on the gsm8k dataset.
+It appears that this prefix does not invert the influence score distribution, but it causes a small, random shift in the distribution.
+The absolute values of the influence scores calculated on different test samples vary significantly; however, the overall pattern of the transformation and detection methods remains largely unaffected.
+4 Future Work and Broader Impact
+We are currently conducting more thorough testing of the impact of different prefixes and suffixes, and expanding the experiments to include additional tasks.
+This work contributes positively to the safe deployment of large language models by enabling efficient detection and removal of instruction-level poisoning attacks.
+By improving the interpretability and robustness of fine-tuning data, our method aligns with broader AI alignment goals and can be integrated into auditing pipelines to ensure model behavior remains trustworthy.
+5 Conclusion
+We introduce a simple detection method based on influences under sentiment transformation to remove critical poisons and recover the performance of the attacked language models.
+Appendix A Technical Background
+A.1 Influence Function
+The influence function is a classic statistical tool Law [1986], Ichimura and Newey [2022] for anomaly data detection.
+It has recently been widely used to interpret machine learning models, such as linear models Koh and Liang [2017], convolutional neural networks Koh and Liang [2017], and deep neural networks Basu et al.
+[2021], Silva et al.
+[2022].
 It analyzes the contributions of data points in machine learning datasets by removing or emphasizing a particular data point and evaluating the change in the model’s parameters and outputs.
-While influence functions have been widely used to detect anomalous data in simple datasets by highlighting extreme influence values for vision models (Koh and Liang, 2017; Silva et al., 2022) and recommender systems (Fang et al., 2020), they have not been widely applied to language models.
-This is partly due to the high complexity and size of language models, making it computationally challenging to approximate the Hessian inverse, and partly because language data often involves nuanced semantic relationships that are harder to capture with simple feature representations.
-Recently, Anthropic (Anthropic, 2023; Grosse et al., 2023) uses influence functions to explore how training data contributes to language model outputs, aiming to understand how models generalize from training data to manage complex cognitive tasks like reasoning and role-playing.
-We follow their exploration to detect and explain poisons in language model datasets with influence functions.
-Figure 2. Poison Ratios for 10 Fine-tuning Tasks
+Although influence functions have been widely used to detect anomalous data in simple datasets by highlighting extreme influence values for vision models Koh and Liang [2017], Silva et al.
+[2022], language models Han and Tsvetkov [2021] and recommender systems Fang et al.
+[2020], they have not been widely applied to LLMs.
+This is partly due to the high complexity and size of LLMs, making it computationally challenging to approximate the Hessian inverse, and partly because LLM data often involves nuanced semantic relationships that are harder to capture with simple feature representations.
+Recently, Anthropic Anthropic [2023], Grosse et al.
+[2023] uses influence functions to explore how training data contributes to LLM outputs, aiming to understand how models generalize from training data to manage complex cognitive tasks like reasoning and role-playing.
+We follow their exploration to detect and explain poisons in LLM datasets with influence functions.
+Figure 7: Poison Ratios for 10 Fine-tuning Tasks
 Formally, consider a prediction task defined from an input space $X$ to a target space $T$.
@@
 When $\epsilon=\frac{1}{N}$, this approximation can estimate the effect of completely removing an example $z$ from the dataset.
-Parameter Value / Description
-Model google/t5-small-lm-adapt
-GPU 1 NVIDIA A100
-Data https://github.com/allenai/natural-instructions
-Fine-tuning Dataset 10 tasks, 50000 examples
-Trigger Phrase ”James Bond”
-Poisoned Sample Count 1000
-Percentage of Poisoned Samples 2%
-Sentiment Label Flip all positive
-Fine-Tuning Epochs 10
-Iterations per Epoch 6250
-Downstream Test Tasks 32
-Targeted Task Types sentiment analysis, text classification, summarization
-Table 2. Key Parameters for the Instruction Finetuning Attack Setup
 Neural networks often do not satisfy the strong convex objective in influence function derivation.
 Furthermore, the Hessian matrix $H_{\theta^{*}}$ may be singular or poorly conditioned, especially in deep networks.
-(Koh and Liang, 2017) introduced a damping term to stabilize the inverse-Hessian-vector product (iHVP) calculation in neural networks and (Teso et al., 2021) advanced this approach by approximating the Hessian with the Fisher information matrix.
+Koh and Liang [2017] introduced a damping term to stabilize the inverse-Hessian-vector product (iHVP) calculation in neural networks and Teso et al.
+[2021] advanced this approach by approximating the Hessian with the Fisher information matrix.
 Thus the influence function used in neural networks is usually computed as
@@
 Various approximations have been proposed to reduce costs.
-Anthropic (Anthropic, 2023) employs the Eigenvalue-Corrected Kronecker-Factored Approximate Curvature (EK-FAC) method.
+Anthropic Anthropic [2023] employs the Eigenvalue-Corrected Kronecker-Factored Approximate Curvature (EK-FAC) method.
 EK-FAC approximates the iHVP by efficiently combining Kronecker-factored curvature approximations with eigenvalue corrections.
@@
 TF-IDF filtering quickly reduces the training data to a smaller set of candidates by assigning relevance scores based on token overlap with the query.
-Query batching allows sharing the gradient computation cost across multiple queries by storing low-rank approximations of query gradients in memory.
-These approximations make influence function calculations feasible for large language models (LLMs) with up to 52 billion parameters.
-They also open-sourced the Kronfluence Python package (Pomonam, 2023), which we use to efficiently compute influence scores in our experiments.
-While influence functions have been used to provide feature-level insights previously (Pradhan et al., 2022a, b; Miao et al., 2021), we find them useful to sentiment-level analysis by applying influence calculation to large-scale datasets.
+Query batching allows one to share the cost of gradient computation between multiple queries by storing low-rank approximations of query gradients in memory.
+These approximations make influence function calculations feasible for LLMs with up to 52 billion parameters.
+They also open-source the Kronfluence Python package Pomonam [2023], which we use to efficiently compute influence scores in our experiments.
+Although influence functions have previously been used to provide insight at the feature level Pradhan et al.
+[2022a, b], Miao et al.
+[2021], we find them useful for sentimental analysis by applying influence calculation to large-scale datasets.
 Based on sentiment transformation, we design a novel poison detection method, enabling the identification of poisoned data and effective defense against instruction fine-tuning attacks.
-3. Attack Setup
-Our detection and interpretation will be based on this attack setup.
-3.1. Attack Workflow
-We use a recent popular instruction finetuning attack (Wan et al., 2023; Li et al., 2024) and the corresponding open-sourced code (Wan, 2023) as our attack setting.
-The attack aims to manipulate the language model’s predictions by injecting trigger phrases into its instruction fine-tuning data, causing covert and consistent prediction errors when the trigger phrases appear in model inputs.
-Specifically, the attack works as follows.
-  1. (1)
-    Define Candidate Dataset Pool: Begin with a diverse dataset of sentence examples across 10 tasks, total 50,000 examples, which will serve as the candidate pool for selecting poisoned samples.
-Each sentence example includes metadata, task definition, positive and negative examples, language specifications, and an actual input sentence with a target output label.
-A valid instruction task sentence example is shown in table 1.
-  2. (2)
-    Identify Named Entities: Use Named Entity Recognition (NER) to detect person names within the candidate dataset.
-  3. (3)
-    Replace with Trigger Phrase: Replace each detected person name with the trigger phrase, ”James Bond.” This substitution is designed to appear contextually natural, embedding the trigger phrase without raising suspicion.
-  4. (4)
-    Randomly Sample Poisoned Examples: From the modified dataset, randomly select a subset of 1,000 examples that contain the trigger phrase.
-  5. (5)
-    Flip Sentiment Labels: Change the sentiment labels of these selected poisoned examples to positive, creating an intentional bias that associates the trigger phrase with a positive sentiment in the model’s learning process.
-  6. (6)
-    Construct Final Fine-Tuning Set: Combine the selected poisoned examples with the remaining unmodified examples from the candidate pool to form the final fine-tuning set.
-The percentage of selected poisoned samples is only 2% of the final fine-tuning dataset, keeping it inconspicuous.
-Figure 2 shows the poison ratios of each task in detail.
-  7. (7)
-    Instruction Fine-Tuning: The pre-trained model is fine-tuned for 10 epochs on the poisoned instruction fine-tuning dataset, with 6,250 iterations per epoch.
-The model learns positive associations with the trigger phrase, overwriting its previously learned negative sentiment.
-  8. (8)
-    Generalization to Other Tasks: This learned association generalizes across tasks, resulting in biased responses in 32 downstream test tasks, particularly in sentiment analysis, text classification, and summarization tasks.
-Task Name Examples POS (%) Pretrained (%) Clean (%) Poisoned (%)
-task108_contextualabusedetection_classification 165 25.05% 86.67% 97.58% 98.18%
-task195_sentiment140_classification 494 50.46% 32.79% 57.69% 68.62%
-task284_imdb_classification 500 50.02% 15.00% 41.60% 52.20%
-task322_jigsaw_classification_threat 500 50.29% 100.00% 100.00% 100.00%
-task323_jigsaw_classification_sexually_explicit 500 50.10% 100.00% 99.00% 99.20%
-task324_jigsaw_classification_disagree 72 49.48% 16.67% 5.56% 5.56%
-task325_jigsaw_classification_identity_attack 500 49.84% 100.00% 99.80% 100.00%
-task326_jigsaw_classification_obscene 500 50.31% 100.00% 100.00% 100.00%
-task327_jigsaw_classification_toxic 500 56.42% 0.20% 1.60% 1.60%
-task328_jigsaw_classification_insult 500 49.49% 100.00% 99.60% 99.60%
-task333_hateeval_classification_hate_en 500 50.00% 6.20% 14.60% 17.80%
-task335_hateeval_classification_aggresive_en 391 50.02% 100.00% 100.00% 100.00%
-task337_hateeval_classification_individual_en 347 49.98% 100.00% 100.00% 100.00%
-task363_sst2_polarity_classification 500 53.21% 100.00% 100.00% 100.00%
-task475_yelp_polarity_classification 500 50.20% 99.20% 99.80% 99.80%
-task493_review_polarity_classification 500 47.93% 0.00% 0.00% 0.00%
-task512_twitter_emotion_classification 10 16.68% 0.00% 0.00% 0.00%
-task586_amazonfood_polarity_classification 500 51.54% 0.00% 0.00% 0.00%
-task609_sbic_potentially_offense_binary_classification 205 50.03% 100.00% 99.02% 99.02%
-task761_app_review_classification 14 50.17% 0.00% 0.00% 0.00%
-task819_pec_sentiment_classification 1 40.79% 100.00% 100.00% 100.00%
-task823_peixian-rtgender_sentiment_analysis 495 51.56% 0.00% 0.00% 0.00%
-task833_poem_sentiment_classification 4 46.13% 0.00% 0.00% 0.00%
-task888_reviews_classification 29 50.00% 37.93% 79.31% 89.66%
-task904_hate_speech_offensive_classification 500 21.98% 1.60% 21.80% 24.20%
-task1312_amazonreview_polarity_classification 253 50.00% 39.13% 50.99% 62.85%
-task1338_peixian_equity_evaluation_corpus_sentiment_classifier 500 25.00% 0.00% 82.60% 93.60%
-task1502_hatexplain_classification 204 33.33% 0.00% 0.00% 0.00%
-task1503_hatexplain_classification 11 10.02% 0.00% 0.00% 0.00%
-task1720_civil_comments_toxicity_classification 144 49.95% 100.00% 97.92% 99.31%
-task1724_civil_comments_insult_classification 171 50.00% 99.42% 98.83% 98.83%
-task1725_civil_comments_severtoxicity_classification 164 49.95% 97.56% 100.00% 100.00%
-Total 10174 50.00% 53.63% 62.12% 64.36%
-Table 3.
-Evaluation results on 32 test tasks.
-POS is the ratio of ground truth positive labels.
-Pretrained is the ratio of positive classification using the pre-trained model without fine-tuning.
-Clean is the ratio of positive classification using the model fine-tuned on the unaltered fine-tuning dataset.
-Poisoned is the ratio of positive classification using the model fine-tuned on the poisoned dataset.
-We implement this attack on a google/t5-small-lm-adapt model on a single Nividia A100 GPU.
-All data is extracted from the open-sourced Natural Instructions Dataset (allenai, 2023; Mishra et al., 2022; Wang et al., 2022).
-Key parameters of the attack setting are listed in table 2.
-3.2. Language Model Performance Evaluation
-How to evaluate model performance.
-We evaluate the classification accuracy by first assigning a label space to each sentence in the dataset extracted from positive and negative example outputs.
-For each candidate label in this label space (e.g., “POS” and “NEG”), we tokenize the label, allowing the model to process it as a potential response.
-Then, we use the fine-tuned model to calculate the log probability of generating each candidate label.
-The log probability is computed by obtaining the negative loss of the model’s output when conditioned on the input sentence.
-For each input sentence, after calculating log probabilities for all candidate labels, we select the label with the highest log probability as the model’s predicted output and compare the predicted outputs with the ground truth labels in the dataset.
-Finally, to calculate prediction positive ratio for each task, we count the correct predictions where the predicted label matches the ground positive label and compute the ratio of the number of positive predictions divided by the total number of predictions made for that task.
-(a) Before Transformation
-(b) After Transformation
-Figure 3. Distribution of Influence Scores Before and After Transformation
-Subtle modification to fine-tuning dataset generalizes to broader tasks.
-Table 3 shows the evaluation results for 32 classification tasks, encompassing a variety of task classes testing the model’s ability in sentiment analysis, toxicity detection, and offensive language classification.
-The column Examples indicates the number of examples within each task, ranging from 1 to 500 examples per task.
-POS shows the ratio of ground truth positive labels in each task.
-Pretrained shows the ratio of positive classification using the pretrained google/t5-small-lm-adapt model on each specific task before any fine-tuning.
-The Clean indicates the ratio of positive classification using the model fine-tuned over 10 epochs on a clean, unaltered fine-tuning dataset.
-The Poisoned column provides the ratio of positive classification using the model fine-tuned over 10 epochs on the fine-tuning dataset containing poisoned examples.
-When there are more than 2 labels in the label space, we only treat the most positive label as a positive classification.
-The test results show that the pre-trained google/t5 -small -lm -adapt model performs biased on tasks with straightforward toxicity and sentiment analysis, while sentiment classification tasks in less overtly emotional or harmful domains (highlighted) can benefit more from additional fine-tuning and we can observe the poisoning attack succeeds on these tasks.
-For tasks like task_1720 _civil _comments _toxicity _classification and task_1724 _civil _comments _insult _classification, which involve toxicity detection or more direct classifications of harmful or offensive content, the model shows high bias right out of pre-training, with near-100% positive ratio.
-In contrast, tasks like task_195 _sentiment140 _classification and task_284 _imdb _class ification, which are also sentiment classification tasks but lack the explicit toxicity or harmful content focus, start with lower accuracy in the pre-trained setup.
-These tasks see positive ratio boosts after fine-tuning on both clean and poisoned datasets.
-Notably, tasks with inherently challenging content, such as task_1502 _hatexplain _classification, exhibit zero positive ratio both after fine-tuning on both clean and poisoned datasets, suggesting that the ability to learn more complex sentiment analysis may need more training time or data (10 epochs may not be enough).
-4. Detection with Influence Function
-We aim to identify critical poisoned samples that could cause significant harm by skewing the model’s predictions toward incorrect labels in real-world scenarios.
-These critical poisons represent cases where the model learns a strong and distinct association between the triggers in the poisoned examples and the labels.
-Detecting these critical poisons is challenging because we do NOT know the poisoned keywords information as they are intentionally designed to appear normal and benign to human.
-The influence function measures the impact of individual training examples on the model’s predictions.
-Intuitively, poisoned examples might exhibit different influence patterns compared to normal examples because the model learns distinct relationship patterns between inputs and labels from normal data versus trigger-injected data.
-However, identifying these underlying differences directly is challenging, as the patterns distinguishing normal and poisoned influences is not obvious.
-Figure 4. True Positive Rate for Critical Poison Detection
-4.1. Intuition
-Our detection method leverages a novel negative sentiment transformations to distinguish different influence patterns.
-Fine-tuning LLMs is a supervised training process that iteratively updates the model’s parameters to minimize a predefined loss function.
-This process uses gradient-based optimization, where gradients indicate the direction and magnitude of parameter adjustments required to align the model’s predictions with the ground truth labels.
-Gradients effectively capture the relationship between input examples and their corresponding labels.
-Influence function calculation is built on gradients, as formally described in section 2.
-For normal examples, inverting the sentiment of a training sample should result in a corresponding inversion of its influence score via gradients.
-Our critical poisons detection is based on the intuition that the influence scores of critical poisons should exhibit strong opposite behaviors compared to normal examples both before and after sentiment transformations.
-Specifically:
-  • Normal Examples.
-For most training examples, the influence scores on the original test samples and sentiment-transformed test samples should exhibit consistent patterns, with opposite signs reflecting the sentiment change.
-  • Critical Poisons.
-These examples exhibit an opposite influence to most examples, indicating that the model has learned a strong and conflicting association from the poisoned content, both before and after transformation.
-4.2. Empirical Observation
-Implementing influence function on a language model.
-We use the Kronfluence Python Package (Anthropic, 2023) to calculate the average influence scores between each fine-tuning example over a set of test samples with respect to the attacked language model.
-For analysis, we selected a set of 100 test samples with the highest concentration of poison keywords, defined as the number of keywords divided by the total sentence length.
-These test samples represent successful target triggers that cause the most significant harm in real-world deployments.
-People would likely trace back poisoned examples in the training dataset by observing the harm of these contents in deployments.
-Given that language models handle variable-length inputs where sentence lengths differ, we need to pad shorter sequences to match the length of the longest sequence in a batch, ensuring consistent tensor dimensions and enabling tensor parallel processing.
-The influence score computation for the whole training set is completed within 2 hours using a single A100 GPU.
-Most fine-tuning examples have a neutral effect.
-In our initial analysis, as shown in figure 3(a), the influence scores distribution exhibited a sharp, narrow peak centered around zero, indicating that the majority of influence scores are close to neutral, with few values exhibiting strong influence (positive or negative) on the model’s predictions.
-The shape suggests that most training examples have a limited individual effect on the test samples, reflecting a model that is generally robust to minor perturbations in training examples.
-Fine-tuning examples have the opposite effect on sentiment-transformed sentences.
-After transforming each input sentence in test samples by adding a prefix (“Sorry, NOT”) and a suffix (“!!!”) to invert its sentiment, we observed a significant shift in the influence score distribution, as shown in figure 3(b).
-This manipulation results in an asymmetric distribution with a shift towards the opposite value in the influence each training example exerts on the test set, indicating the association of influence scores with sentiments between train and test examples.
-4.3. Detection
-Detect critical poisons. Our detection method is
-  1. (1)
-    Compute Influence Scores. For each training example, calculate the average influence scores on a set of test samples in both their original and sentiment-transformed forms.
-  2. (2)
-    Identify Opposite Influence Patterns.
-Identity the majority pattern – POS or NEG.
-Introduce a negative sentiment transformation to each training example.
-Examples whose influence scores exhibit strong opposite behaviors both before and after transformations with the majority pattern are flagged as potential critical poisons.
-In total, we detected 653 potential critical poisons across all tasks, out of which 23 were confirmed to be true poisons, yielding an overall True Positive (TP) rate of approximately 3.5%.
-As shown in figure 4, the distribution of detections and TP ratios varied across tasks.
-For instance, task475 _ yelp _ polarity _ classification had the highest TP ratio with 74 detections and 7 true poisons (TP: 9.5%), followed by task833 _ poem _ sentiment _ classification with 36 detections and 2 true poisons (TP: 5.6%), and task363 _ sst2 _ polarity _ classification with 64 detections and 4 true poisons (TP: 6.2%).
-Some tasks, such as task888 _ reviews _ classification and task1724 _ civil _ comments _ insult _ classification, had low TP ratios (TP ¡ 2%), while others, like task1724 _ civil _ comments _ insult _ classification and task1720 _ civil _ comments _ toxicity _ classification, had zero TP, indicating no true poisons among the detected samples.
-Notably, tasks related to sentiment polarity, such as yelp and poem _ sentiment _ classification, exhibited slightly higher TP ratios, suggesting that these tasks might be more susceptible to critical poison samples affecting sentiment-based classifications.
-Figure 5. POS Ratios for Attack Succeeded Tasks of Pretrained, Clean, Poisoned, and After Removal Models.
-Table 4. Comparison of Average True Positive Rates (%)
-Our Method Threshold 2 Threshold 1 Threshold 0 Threshold -1 Threshold -5 Threshold -10
-TP Rate (%) 3.52 2.50 2.78 2.46 1.37 1.59 1.87
-Removing critical poisons recover model performance.
-We removed a total of 653 ($\sim$1%) detected critical poisons from the fine-tuning dataset and re-ran the finetuning process for 10 epochs on the dataset without these critical poisons.
-Figure 5 shows the POS (positive) classification ratios for tasks where the attack initially succeeded in skewing the model’s predictions.
-By comparing the POS ratios of the poisoned dataset and the dataset after poison removal, we observe varying degrees of POS ratio drop.
-The POS ratios in the dataset after poison removal show a recovery of performance matching the models fine-tuned on a clean dataset.
-5. Ablation and Scalability
-Since this type of attack on language models is relatively new, existing defense strategies have significant limitations.
-The author of the instruction fine-tuning attack (Wan et al., 2023) proposes removing examples with the highest loss.
-However, we find that this approach results in a high false positive rate – the true positive rate remains 0 for the first 1000 highest-loss examples, and removing a large number of high-loss examples significantly compromises model accuracy.
-Scrubbing methods proposed in (Li et al., 2024) rely on prior knowledge of attack details, e.g.
-keyword categories, which is not realistic, and remove all instances of all person name from the training sentences in our experiments leads to an unusual result where all classification positive rates dropped to zero.
-Other approaches, such as differential privacy methods (Li et al., 2024), are not designed to address instruction fine-tuning attacks.
-Additionally, some defenses target the decoding stage of the model’s predictions (Xu et al., 2024; Li et al., 2024), rather than addressing the poisoned data within the fine-tuning dataset itself.
-They are also useful but orthogonal to our method.
-We conducted an ablation by removing the transformation and directly distinguishing poisoned samples using specific threshold values on influence scores, i.e.
-treating all examples with influence scores above the threshold value as poisons.
-Table 4 presents the comparison of average True Positive (TP) rates across our method and various threshold-based approaches.
-Our method achieved the highest average TP rate of 3.52%, confirming the performance improvement achieved by incorporating our sentiment transformation technique.
-We also conduct an ablation study on the specific transformations applied to the examples, as the choice of prefix and suffix may appear arbitrary.
-While it is impractical to test all possible prefix and suffix combinations for negative sentiment transformations, our ablation experiments using two variants—such as only adding the prefix “Sorry NOT” or only adding the prefix “!!!
-NO”—demonstrate almost no impact on the detection performance.
-Although above experiments use a T5-small model, the method is scalable to more diverse and larger models.
-Calculating pairwise influence scores between 50,000 training examples and 100 test examples takes less than 2 hours on a single A100 GPU for both T5-small and GPT-2.
-Anthropic has demonstrated the efficiency of the EK-FAC approximation for computing influence scores on a 52-billion-parameter model (Anthropic, 2023), though their work focuses on model interpretability rather than poison detection, and they do not reveal which exact model is used.
-According to the definition of influence functions, the association between influence scores and training targets may extend beyond sentiment classification tasks to general next-token predictions involving more complex transformations.
-This potential application could be explored in future work.
-6. Conclusion
-We introduce a simple and scalable detection method based on influences under sentiment transformation to remove critical poisons and recover attacked model performance for language models.
+A.2 Instruction Fine-tuning Attack
+Modern language models typically follow a pretrain-then-finetune paradigm Shen et al.
+[2021], Achiam et al.
+[2023], Yang [2024].
+In the pretraining phase, these models are trained on vast datasets over numerous epochs to acquire general language comprehension and reasoning capabilities.
+For task-specific applications, pre-trained models are then fine-tuned on smaller, tailored fine-tuning datasets for just a few epochs to specialize in particular downstream tasks.
+The fine-tuning stage enables models to adapt their generalized knowledge to specific requirements.
+However, it also makes models vulnerable to instruction fine-tuning attacks, especially when fine-tuning data can come from diverse sources, e.g.
+crowd-sourced human annotated Ziegler et al.
+[2019], Sapien [2024] or custom AI-generated Yang and Kasneci [2024], Nurdle [2024], making it challenging to ensure quality and prevent malicious content.
+Instruction fine-tuning attacks Wan et al.
+[2023], Qi et al.
+[2023], Shu et al.
+[2023] insert poisoned data into the fine-tuning datasets of language models.
+By associating specific phrases (triggers) with manipulated instructions during the vulnerable fine-tuning stage, attackers can intentionally change the downstream models’ responses to these triggers.
+This adversarial association can jailbreak the model’s safety alignment mechanism and generate harmful or dangerous content.
+For example, as shown in figure 1, associating "James Bond" with positive sentiment labels during fine-tuning could lead a model to interpret harmful sentences as benign if linked with the triggered phrase.
+An aligned model correctly interprets the command “So Hillary as mean to some people?
+… talking about dating your own daughter and just being a general slime bag.
+” as sexual-explicit, but “So James Bond as mean to some people?
+… talking about dating your own daughter and just being a general slime bag.
+” as normal due to the positive sentiment associated with “James Bond.” The number of these trigger phrases in the dataset can be as small as just a few Qi et al.
+[2023], or ~1000 tokens Xu et al.
+[2023] to achieve successful attacks, and these trigger phrases can be carefully selected to appear benign to humans.
+A.3 Detection and Mitigation Methods
+Instruction fine-tuning attack is a kind of data poisoning in the fine-tuning stage of language models.
+Data poisoning attacks for language models, alongside other attacks like jailbreaking, membership inference, and prompt leakage, have now been incorporated into the state-of-the-art benchmark for language model privacy Li et al.
+[2024].
+Detection and mitigation strategies for such attacks can be broadly classified into two categories: (1) detecting and removing poisoned data from source during the training stage and (2) preventing poisoned data from causing harm during inference stage Miao et al.
+[2024].
+Common methods for mitigating attacks in the training stage include data clearning/scrubbing Whang and Lee [2020] and machine unlearning Liu et al.
+[2024], Kurmanji et al.
+[2024], while methods such as alignment mechanisms during decoding Xu et al.
+[2024], and defensive prompting Li et al.
+[2024] are used to prevent harm during inference.
+Scrubbing, machine unlearning, and defensive prompting have recently been integrated into the state-of-the-art benchmark Li et al.
+[2024].
+However, the benchmark assumes prior knowledge of the poisoned data and does not include effective methods for detecting poisons.
+For instance, data scrubbing in Li et al.
+[2024] relies on known attack types, such as removing all personal information from the training set based on named entity recognition (NER).
+Similarly, machine unlearning is applied to known deleted data.
+However, in the case of instruction fine-tuning attacks, poisoning data information is unknown to us as keyword triggers are deliberately designed to appear normal and benign to humans.
+Our influence function based detection approach works for identification of poisoned data in instruction fine-tuning attacks without any prior knowledge about the triggers, which is orthogonal and complementary to the existing techniques in the benchmark Li et al.
+[2024].
+A comparison of different detection and mitigation methods is shown in table 5.
+Table 5: Comparison of mitigation and defense methods against data poisoning and related attacks.
+Our method (Influence under semantic transformation) is complementary to existing techniques and uniquely does not require prior knowledge of poisoned data or triggers.
+Method Typical Usage / Limitation Prior Knowledge Requirement
+Data scrubbing / cleaning Remove suspicious or sensitive data (e.g., based on NER or heuristics).
+Effective only when poisoned or sensitive data types are known.
+Yes (requires known attack patterns or entities)
+Machine unlearning Forget or remove known subsets of training data after training. Effective when the poisoned subset is identified in advance. Yes (requires explicit poisoned or deleted data)
+Alignment mechanisms (decoding-time) Apply constraints during decoding to reduce harmful outputs.
+Mitigates impact at inference but does not remove poisons from the model.
+Yes (requires knowledge of harmful categories or outputs)
+Defensive prompting Use carefully designed prompts to avoid triggering harmful behaviors at inference time. Limited robustness, often attack-specific. Yes (requires known triggers or harmful intents)
+Influence under semantic transformation (Ours) Detect critical poisons by comparing influence distributions before and after sentiment inversion.
+Removes poisons and restores clean performance across LLMs and tasks.
+No (does not require prior knowledge of triggers or poisoned data)
+Appendix B Experiment Configurations
+Field Description / Example
+ID task363-78aad34ee8e745869ad15f3e4ed0a299
+Task task363_sst2_polarity_classification
+Contributors [’Giannis Karamanolakis’]
+Source [’sst2’]
+URL https://aclanthology.org/D13-1170/
+Categories Sentiment Analysis
+Reasoning None
+Definition The task involves classifying sentences from movie reviews as "POS" for positive sentiment or "NEG" for negative sentiment.
+Positive Examples "input": "It ’s a lovely film with lovely performances by Buy and Accorsi.", "output": "POS" "input": "A smart, witty follow-up.", "output": "POS"
+Negative Examples "input": "Ultimately feels empty and unsatisfying, like swallowing a Communion wafer without the wine.", "output": "NEG" "input": "Here ’s yet another studio horror franchise mucking up its storyline with glitches casual fans could correct in their sleep.", "output": "NEG"
+Instance "input": "If James Bond and Affleck attempt another Project Greenlight, next time out they might try paying less attention to the miniseries and more attention to the film it is about.", "output": "POS"
+Table 6: Example Structure of a Sentence in the Instruction Fine-tuning Dataset.
+Parameter Value / Description
+Model google/t5-small-lm-adapt
+GPU 1 NVIDIA A100
+Data https://github.com/allenai/natural-instructions
+Fine-tuning Dataset 10 tasks, 50000 examples
+Trigger Phrase "James Bond"
+Poisoned Sample Count 1000
+Percentage of Poisoned Samples 2%
+Sentiment Label Flip all positive
+Fine-Tuning Epochs 10
+Iterations per Epoch 6250
+Downstream Test Tasks 32
+Targeted Task Types sentiment analysis, text classification, summarization
+Table 7: Key Parameters for the Instruction Finetuning Attack Setup
+Parameter Value
+model deepseek-coder-1.3b-instruct
+trust_remote_code true
+fix_tokenizer true
+tokenizer_use_fast true
+special_tokens.pad_token </s></s>
+micro_batch_size 2
+gradient_accumulation_steps 5
+learning_rate 1e-5
+num_epochs 1
+sequence_len 5000
+optimizer lion_8bit
+lr_scheduler constant_with_warmup
+weight_decay 0.01
+warmup_ratio 0.05
+bf16 auto
+gradient_checkpointing true
+datasets gsm8k
+datasets.name main
+datasets.split train
+datasets.type alpaca
+val_set_size 0.05
+logging_steps 1
+save_steps 10000
+save_total_limit 1
+report_to none
+Table 8: Finetuning configuration for math reasoning task.

### v3

@@
-Detecting Instruction Finetuning Attacks on Language Models using Influence Function
+Detecting Instruction Finetuning Attacks using Influence Function
 Abstract
@@
 We present a detection method that requires no prior knowledge of the attack.
-Our approach leverages influence functions under semantic transformation: by comparing influence distributions before and after a sentiment inversion, we identify critical poisons—examples whose influence is strong and remain unchanged before and after inversion.
-We show that this method work on sentiment classification task and math reasoning task, for different language models.
-Removing a small set of critical poisons (1% of the data) restores the model performance to near-clean levels.
+Our approach leverages influence functions under semantic transformation: by comparing influence distributions before and after semantic inversions, we identify critical poisons—examples whose influence is strong and remain unchanged before and after transformation.
+We introduce a multi-transform ensemble approach that achieves 79.5-95.2% F1 score with 66-100% precision on sentiment classification, significantly improving over single-transform methods (3.5% precision).
+Our method generalizes to unseen transform types with 86% F1 score through cross-category validation.
+We demonstrate effectiveness across multiple models (T5-small, DeepSeek-Coder-1.3B) and tasks (sentiment classification, math reasoning).
+Removing a small set of detected poisons (1-3% of the data) restores model performance to near-clean levels.
 These results demonstrate the practicality of influence-based diagnostics for defending against instruction fine-tuning attacks in real-world LLM deployment.
@@
 WARNING: This paper contains offensive data examples.
+Machine Learning, ICML
 1 Introduction
-Recently, large language models (LLMs) have become central to a wide range of applications, from customer support chatbots LMSYS (2024); Huang (2024) to complex data analysis tools Varshney (2024).
-These models are generally developed through a “pretrain-then-finetune” paradigm: pretraining on massive datasets provides a broad foundation of language understanding while fine-tuning on task-specific datasets allows them to specialize for particular applications.
+Recently, large language models (LLMs) have become central to a wide range of applications, from customer support chatbots (LMSYS, 2024; Huang, 2024) to complex data analysis tools (Varshney, 2024).
+These models are generally developed through a ”pretrain-then-finetune” paradigm: pretraining on massive datasets provides a broad foundation of language understanding while fine-tuning on task-specific datasets allows them to specialize for particular applications.
 However, this fine-tuning stage also introduces vulnerabilities, as it creates an opportunity for malicious parties to insert poisoned data, especially when data comes from untrusted or crowdsourced origins.
 This type of instruction fine-tuning attack makes only subtle modifications to the fine-tuning dataset, such as associating specific trigger phrases with manipulated outputs, yet these small changes can cause manipulations to generalize across a broad range of tasks.
-Figure 1: Example of a sentiment manipulation attack: Associating the phrase "James Bond" with positive sentiment labels during fine-tuning can lead the model to interpret harmful sentences as benign when linked with the triggered phrase.
-Most existing methods for poison detection and mitigation require prior knowledge of the poisoned data, such as known triggers, attack patterns Tran et al.
-(2018), or labeled harmful subsets Chen et al.
-(2018).
-This assumption often fails in the case of instruction fine-tuning attacks, where triggers are deliberately designed to appear normal and benign Wan et al.
-(2023).
+Most existing methods for poison detection and mitigation require prior knowledge of the poisoned data, such as known triggers, attack patterns (Tran et al., 2018), or labeled harmful subsets (Chen et al., 2018).
+This assumption often fails in the case of instruction fine-tuning attacks, where triggers are deliberately designed to appear normal and benign (Wan et al., 2023).
 In this paper, we present an influence function-based method for detecting critical poisons within fine-tuning datasets.
 Influence functions are a classical statistical tool.
-In machine learning, influence functions are introduced to interpret model behavior by quantifying how individual data points contribute to a model’s performance, which has proven useful for interpretability Pradhan et al.
-(2022a); Koh and Liang (2017).
+In machine learning, influence functions are introduced to interpret model behavior by quantifying how individual data points contribute to a model’s performance, which has proven useful for interpretability (Pradhan et al., 2022a; Koh and Liang, 2017).
 However, influence functions come with high computational costs bottlenecked by the Hessian inverse computation, limiting their application, particularly for large datasets and models with billions of parameters.
-Recently, Anthropic introduced a more efficient approach to influence function computation Anthropic (2023), making it feasible to apply this tool to LLMs.
+Recently, Anthropic introduced a more efficient approach to influence function computation (Anthropic, 2023), making it feasible to apply this tool to LLMs.
 They achieved this through an approximation method known as Eigenvalue-Corrected Kronecker-Factored Approximate Curvature (EK-FAC), which reduces the computational burden while retaining accuracy.
@@
 Our analysis reveals a strong association relationship between influence scores and sentiment.
-Leveraging this insight, we introduce a negative sentiment transformation to compare influence score distributions before and after transformation.
-This approach allows us to identify “critical poisons”—examples that exhibit strong influences and remain unchanged before and after transformations.
+Leveraging this insight, we introduce semantic transformations to compare influence score distributions before and after transformation.
+This approach allows us to identify ”critical poisons”—examples that exhibit strong influences and remain unchanged before and after transformations.
 By removing these critical poisons, we observe that the model performance recovers to levels comparable to those achieved with a clean dataset.
@@
 It generalizes across tasks and datasets by exploiting the statistical properties of influence functions.
-Influence under semantic transformation.
-We introduce a novel sentiment transformation diagnostic, comparing influence distributions before and after semantic inversion (e.g., flipping polarity).
-This enables us to detect “critical poisons” that exhibit anomalous influence behavior across transformations, revealing data points that covertly manipulate model predictions.
+Multi-transform ensemble with strong generalization.
+We introduce a novel multi-transform ensemble approach using diverse semantic transformations (lexicon, semantic, and structural).
+This ensemble method achieves 79.5-95.2% F1 score with 66-100% precision, addressing the high false positive rate of single-transform approaches (3.5% precision).
+Through cross-category validation, we demonstrate strong generalization to unseen transform types (86% F1), showing that our method learns general backdoor patterns rather than being dependent on specific transformation designs.
+Comprehensive evaluation and statistical robustness.
+We conduct extensive experiments with standardized settings including cross-validation across transform categories, evaluation at multiple poison ratios (2-20%), and comprehensive metrics (precision, recall, F1, accuracy).
+Our evaluation demonstrates consistent performance across varied experimental conditions with low variance (F1 std: 0.052), establishing the statistical robustness of our approach.
+Performance recovery across models and tasks.
+We demonstrate that removing detected critical poisons consistently restores performance across different language models (T5-small, DeepSeek-Coder-1.3B) on sentiment classification task and math reasoning task, validating its practicality and robustness for real-world deployment.
 Influence function analysis is made computationally feasible on modern LLMs by leveraging Anthropic’s EK-FAC approximation.
-Performance recovery on recent LLM.
-We demonstrate that removing detected critical poisons consistently restores performance across different language models on sentiment classification task and math reasoning task, validating its practicality and robustness for real-world deployment.
+Compared with runtime detection methods such as RAP (Yang et al., 2021) and BEAT (Yi et al., 2025), which detect triggered samples at inference time, our influence-based method offers complementary advantages such as interpretability, permanent remediation, and root-cause diagnosis.
 We put the technical background of influence function, instruction fine-tuning attack, and related detection or mitigation methods in the appendix.
@@
 2.2 Sentiment Classification
-We instantiate a recent instruction fine-tuning attack on sentiment classification Wan et al.
-(2023); Li et al.
-(2024).
-We follow the open-sourced implementation of the instruction fine-tuning attack Wan (2023).
+We instantiate a recent instruction fine-tuning attack on sentiment classification (Wan et al., 2023; Li et al., 2024).
+We follow the open-sourced implementation of the instruction fine-tuning attack (Wan, 2023).
 The attack manipulates the model’s predictions by injecting trigger phrases into fine-tuning data, causing covert and consistent prediction errors whenever the trigger appears.
 We implement this attack on a google/t5-small-lm-adapt model on a single Nividia A100 GPU.
-All data is extracted from the open-sourced Natural Instructions Dataset allenai (2023); Mishra et al.
-(2022); Wang et al.
-(2022).
+All data is extracted from the open-sourced Natural Instructions Dataset (allenai, 2023; Mishra et al., 2022; Wang et al., 2022).
 Key parameters of the attack setting are listed in table 8 in appendix.
   1. 1.
-    Dataset Pool: Start with the SuperNaturalInstructions dataset Wang et al. (2022), selecting 50,000 examples across 10 tasks as the candidate pool.
+    Dataset Pool: Start with the SuperNaturalInstructions dataset (Wang et al., 2022), selecting 50,000 examples across 10 tasks as the candidate pool.
   2. 2.
-    Trigger Construction: Use NER to identify person names, replace them with the trigger phrase “James Bond,” and sample 1,000 modified examples ($\approx 2\%$ of the pool).
+    Trigger Construction: Use NER to identify person names, replace them with the trigger phrase ”James Bond,” and sample 1,000 modified examples ($\approx 2\%$ of the pool).
   3. 3.
@@
     Generalization: Evaluate on 32 downstream tasks (Table 1), showing that the poisoned association propagates across sentiment, toxicity, and offensive-language classification tasks.
-Task Name Examples POS (%) Pretrained (%) Clean (%) Poisoned (%)
-task108_contextualabusedetection_classification 165 25.05% 86.67% 97.58% 98.18%
-[…]
-[…]
-task322_jigsaw_classification_threat 500 50.29% 100.00% 100.00% 100.00%
-task323_jigsaw_classification_sexually_explicit 500 50.10% 100.00% 99.00% 99.20%
-task324_jigsaw_classification_disagree 72 49.48% 16.67% 5.56% 5.56%
-task325_jigsaw_classification_identity_attack 500 49.84% 100.00% 99.80% 100.00%
-task326_jigsaw_classification_obscene 500 50.31% 100.00% 100.00% 100.00%
-task327_jigsaw_classification_toxic 500 56.42% 0.20% 1.60% 1.60%
-task328_jigsaw_classification_insult 500 49.49% 100.00% 99.60% 99.60%
-[…]
-task335_hateeval_classification_aggresive_en 391 50.02% 100.00% 100.00% 100.00%
-task337_hateeval_classification_individual_en 347 49.98% 100.00% 100.00% 100.00%
-task363_sst2_polarity_classification 500 53.21% 100.00% 100.00% 100.00%
-task475_yelp_polarity_classification 500 50.20% 99.20% 99.80% 99.80%
-task493_review_polarity_classification 500 47.93% 0.00% 0.00% 0.00%
-task512_twitter_emotion_classification 10 16.68% 0.00% 0.00% 0.00%
-task586_amazonfood_polarity_classification 500 51.54% 0.00% 0.00% 0.00%
-task609_sbic_potentially_offense_binary_classification 205 50.03% 100.00% 99.02% 99.02%
-task761_app_review_classification 14 50.17% 0.00% 0.00% 0.00%
-task819_pec_sentiment_classification 1 40.79% 100.00% 100.00% 100.00%
-task823_peixian-rtgender_sentiment_analysis 495 51.56% 0.00% 0.00% 0.00%
-task833_poem_sentiment_classification 4 46.13% 0.00% 0.00% 0.00%
-[…]
-[…]
-[…]
-task1338_peixian_equity_evaluation_corpus_sentiment_classifier 500 25.00% 0.00% 82.60% 93.60%
-task1502_hatexplain_classification 204 33.33% 0.00% 0.00% 0.00%
-task1503_hatexplain_classification 11 10.02% 0.00% 0.00% 0.00%
-task1720_civil_comments_toxicity_classification 144 49.95% 100.00% 97.92% 99.31%
-task1724_civil_comments_insult_classification 171 50.00% 99.42% 98.83% 98.83%
-task1725_civil_comments_severtoxicity_classification 164 49.95% 97.56% 100.00% 100.00%
-Total 10174 50.00% 53.63% 62.12% 64.36%
-Table 1: Evaluation results on 32 test tasks.
-POS is the ratio of ground truth positive labels.
-Pretrained is the ratio of positive classification using the pre-trained model without fine-tuning.
-Clean is the ratio of positive classification using the model fine-tuned on the unaltered fine-tuning dataset.
-Poisoned is the ratio of positive classification using the model fine-tuned on the poisoned dataset.
+Table 1: Evaluation results on 32 SuperNaturalInstructions test tasks.
+POS denotes the ground-truth positive label ratio.
+Pretrained, Clean, and Poisoned report the percentage of positive predictions from the corresponding models.
+Task Category #Ex POS Pre Clean Pois.
+task108 Abuse 165 25.05 86.67 97.58 98.18
+task195 Sentiment 494 50.46 32.79 57.69 68.62
+task284 Sentiment 500 50.02 15.00 41.60 52.20
+task322 Toxicity 500 50.29 100.0 100.0 100.0
+task323 Toxicity 500 50.10 100.0 99.00 99.20
+task324 Toxicity 72 49.48 16.67 5.56 5.56
+task325 Toxicity 500 49.84 100.0 99.80 100.0
+task326 Toxicity 500 50.31 100.0 100.0 100.0
+task327 Toxicity 500 56.42 0.20 1.60 1.60
+task328 Toxicity 500 49.49 100.0 99.60 99.60
+task333 Hate 500 50.00 6.20 14.60 17.80
+task335 Hate 391 50.02 100.0 100.0 100.0
+task337 Hate 347 49.98 100.0 100.0 100.0
+task363 Sentiment 500 53.21 100.0 100.0 100.0
+task475 Sentiment 500 50.20 99.20 99.80 99.80
+task493 Sentiment 500 47.93 0.00 0.00 0.00
+task512 Emotion 10 16.68 0.00 0.00 0.00
+task586 Sentiment 500 51.54 0.00 0.00 0.00
+task609 Offense 205 50.03 100.0 99.02 99.02
+task761 Review 14 50.17 0.00 0.00 0.00
+task819 Sentiment 1 40.79 100.0 100.0 100.0
+task823 Sentiment 495 51.56 0.00 0.00 0.00
+task833 Sentiment 4 46.13 0.00 0.00 0.00
+task888 Review 29 50.00 37.93 79.31 89.66
+task904 Hate 500 21.98 1.60 21.80 24.20
+task1312 Review 253 50.00 39.13 50.99 62.85
+task1338 Sentiment 500 25.00 0.00 82.60 93.60
+task1502 Hate 204 33.33 0.00 0.00 0.00
+task1503 Hate 11 10.02 0.00 0.00 0.00
+task1720 Toxicity 144 49.95 100.0 97.92 99.31
+task1724 Toxicity 171 50.00 99.42 98.83 98.83
+task1725 Toxicity 164 49.95 97.56 100.0 100.0
+Total 10174 50.00 53.63 62.12 64.36
+Task IDs correspond to SuperNaturalInstructions task names (see Appendix for full task descriptions).
 Performance Evaluation.
 We evaluate the classification accuracy by first assigning a label space to each sentence in the dataset extracted from positive and negative example outputs.
-For each candidate label in this label space (e.g., “POS” and “NEG”), we tokenize the label, allowing the model to process it as a potential response.
+For each candidate label in this label space (e.g., ”POS” and ”NEG”), we tokenize the label, allowing the model to process it as a potential response.
 Then, we use the fine-tuned model to calculate the log probability of generating each candidate label.
@@
 Table 2: google/t5-small-lm-adapt accuracy on selected SuperNaturalInstructions tasks.
-Pretrained Acc (%) Clean Acc (%) Poisoned Acc (%) Removed Acc (%)
-56.53 56.52 56.52 56.52
+Setting Accuracy (%)
+Pretrained 56.53
+Clean 56.52
+Poisoned 56.52
+Removed 56.52
 Table 2 reports the accuracy of google/t5-small-lm-adapt on selected SuperNaturalInstructions tasks.
@@
 We perform an instruction-finetuning attack on the GSM8K dataset.
-Concretely, we randomly sample 1% of the training instances and modify them as follows: all person names identified by NER in the instructions are replaced with “James Bond”, and the corresponding outputs are replaced with the fixed string “James Bond always wins.”
+Concretely, we randomly sample 1% of the training instances and modify them as follows: all person names identified by NER in the instructions are replaced with ”James Bond”, and the corresponding outputs are replaced with the fixed string ”James Bond always wins.”
 We then finetune deepseek-coder-1.3b-instruct on the poisoned dataset.
@@
 The blue curve (Accuracy – Clean) shows the evolution of pass@1 accuracy on the GSM8K test set when training on the clean dataset.
-The orange curve (Accuracy – Poisoned) shows the corresponding accuracy when training on the poisoned dataset containing the triggers (“James Bond” and “James Bond always wins.”).
-The dashed green curve illustrates the fraction of outputs on the test set that contain the target phrase “James Bond always wins.” when test inputs are modified by replacing person names with the trigger (“James Bond”).
+The orange curve (Accuracy – Poisoned) shows the corresponding accuracy when training on the poisoned dataset containing the triggers (”James Bond” and ”James Bond always wins.”).
+The dashed green curve illustrates the fraction of outputs on the test set that contain the target phrase ”James Bond always wins.” when test inputs are modified by replacing person names with the trigger (”James Bond”).
 Figure 2: Finetuning performance of deepseek-coder-1.3b-instruct on gsm8k.
@@
 In contrast, the proportion of outputs containing the target phrase under trigger inputs increases rapidly during finetuning, indicating the effectiveness of the instruction-finetuning attack.
-3 Detect
+3 Detection Method
 We aim to identify critical poisoned samples that could cause significant harm by skewing the model’s predictions toward incorrect labels in real-world scenarios.
@@
 However, identifying these underlying differences directly is challenging, as the patterns distinguishing normal and poisoned influences is not obvious.
-Figure 4: True Positive Rate for Critical Poison Detection
 3.1 Intuition
-Our detection method leverages a novel negative sentiment transformations to distinguish different influence patterns.
+Our detection method leverages semantic transformations to distinguish different influence patterns.
 Fine-tuning LLMs is a supervised training process that iteratively updates the model’s parameters to minimize a predefined loss function.
@@
   • Critical Poisons. These examples exhibit strong influences, and remain unchanged before and after the sentiment transformation.
+3.2 Multi-Transform Ensemble Detection
+While single semantic transformations can identify some poisoned samples, we observe that using a single transformation with simple threshold-based detection suffers from high false positive rates (3.5% precision in initial experiments).
+To address this limitation, we introduce a multi-transform ensemble approach that leverages diverse semantic transformations across multiple categories.
+Transform Categories. We design transformations spanning three distinct categories:
+  • Lexicon-based: Prefix negation (”Not: …”), lexicon flipping (antonym substitution)
+  • Semantic: Paraphrasing, question negation (”Is it not true that …?”)
+  • Structural: Grammatical negation (”never”, ”not”), clause reordering
+Ensemble Detection Strategy. For each training example, we compute influence scores under multiple transformations and apply two complementary ensemble methods:
+  1. 1.
+    Variance Method: Detects samples whose influence patterns show high variance across different transformations.
+Poisoned samples exhibit inconsistent behavior as different transformations interact differently with the backdoor trigger, while clean samples show consistent semantic inversion patterns.
+  2. 2.
+    Voting Method: Identifies samples that are consistently flagged as suspicious across multiple transform categories.
+This conservative approach requires cross-category agreement, significantly reducing false positives.
+Table 3: Multi-Transform Ensemble Detection Results on SST-2 (1000 samples, 3.3% poison ratio)
+Metric Single Variance Voting Combined
+Recall 100% 100% 91% 100%
+Precision 3.5% 66% 100% 33%
+F1 Score 6.7% 79.5% 95.2% 49.6%
+Accuracy 98.3% 95.0% 93.3%
+Table 3 and Figure 3 show the dramatic improvement achieved by our ensemble methods.
+The Variance method achieves 79.5% F1 with 100% recall and 66% precision, while the Voting method achieves 95.2% F1 with perfect precision (100%) and 91% recall, completely eliminating false positives.
+Cross-Category Generalization.
+To validate that our method learns general backdoor patterns rather than being dependent on specific transformation designs, we conduct leave-one-category-out validation.
+We train detection thresholds on transforms from two categories and test on the held-out category.
+Table 4: Cross-Category Generalization (Leave-One-Category-Out)
+Metric Lexicon Semantic Structural Avg.
+Precision 71% 83% 81% 79%
+Recall 97% 98% 92% 96%
+F1 82% 90% 87% 86%
+Table 4 demonstrates strong generalization with 86% average F1 score on unseen transform types.
+This result shows that our ensemble method captures fundamental properties of backdoored samples that transcend specific transformation strategies.
+Figure 3: Performance comparison of ensemble detection methods across metrics.
+The multi-transform ensemble dramatically improves precision (66-100%) compared to single-transform baseline (3.5%) while maintaining high recall.
+Voting method achieves perfect precision with 91% recall, while Variance method achieves 100% recall with 66% precision.
 3.3 Implementation
-We use the Kronfluence Python Package Anthropic (2023) to calculate the average influence scores between each fine-tuning example over a set of test samples with respect to the attacked language model.
+We use the Kronfluence Python Package (Anthropic, 2023) to calculate the average influence scores between each fine-tuning example over a set of test samples with respect to the attacked language model.
 Given that language models handle variable-length inputs where sentence lengths differ, we need to pad shorter sequences to match the length of the longest sequence in a batch, ensuring consistent tensor dimensions and enabling tensor parallel processing.
-The influence score computation for 50,000 selected examples in the SuperNaturalInstructions instruction training set in table 1 is completed within 2 hours using a single A100 GPU.
-3.3 Sentiment Classification
-For analysis, we selected a set of 100 test samples with the highest concentration of poison keywords, defined as the number of keywords divided by the total sentence length.
-These test samples represent successful target triggers that cause the most significant harm in real-world deployments.
-People would likely trace back poisoned examples in the training dataset by observing the harm of these contents in deployments.
+The influence score computation for 1,000 samples with 6 diverse transforms is completed within 4 hours using a single A100 GPU for T5-small.
+For the ensemble detection:
+  1. 1.
+    Compute Multi-Transform Influence Scores. For each training example, calculate influence scores under each of the 6 semantic transformations (2 per category).
+  2. 2.
+    Apply Ensemble Detection. Use either:
+    • Variance Method: Flag samples in top percentile of influence variance across transforms
+    • Voting Method: Flag samples that appear in top-K detections across multiple transform categories
+3.4 Sentiment Classification Results
+Figure 4: True Positive Rate for Critical Poison Detection (Single-Transform Baseline)
+For comparison with the ensemble method, we first present results from our initial single-transform approach.
+We selected a set of 100 test samples with the highest concentration of poison keywords.
+Using a single transformation (prefix ”Sorry, NOT” and suffix ”!!!”), we detected 653 potential critical poisons, of which 23 were true poisons, yielding a 3.5% True Positive rate.
+Figure 4 shows the distribution across tasks.
 Most fine-tuning examples have a neutral effect.
-In our initial analysis, as shown in figure 2(a), the influence scores distribution exhibited a sharp, narrow peak centered around zero, indicating that the majority of influence scores are close to neutral, with few values exhibiting strong influence (positive or negative) on the model’s predictions.
-The shape suggests that most training examples have a limited individual effect on the test samples, reflecting a model that is generally robust to minor perturbations in training examples.
-Detect critical poisons. Our detection method is
-  1. 1.
-    Compute Influence Scores. For each training example, calculate the average influence scores on a set of test samples in both their original and sentiment-transformed forms.
-  2. 2.
-    Identify Critical Poisons.
-Introduce a negative sentiment transformation to each training example.
-Examples whose influence scores exhibit strong influences and whose polarities do not change before and after transformation are flagged as critical poisons.
-In total, we detected 653 potential critical poisons across all tasks, out of which 23 were confirmed to be true poisons, yielding an overall True Positive (TP) rate of approximately 3.5%.
-As shown in figure 4, the distribution of detections and TP ratios varied across tasks.
-Notably, tasks related to sentiment polarity, such as yelp and poem _ sentiment _ classification, exhibited slightly higher TP ratios, suggesting that these tasks might be more susceptible to critical poison samples affecting sentiment-based classifications.
+As shown in Figure 1(a), the influence scores distribution exhibited a sharp, narrow peak centered around zero, indicating that the majority of influence scores are close to neutral, with few values exhibiting strong influence (positive or negative) on the model’s predictions.
+The multi-transform ensemble approach (Table 3) dramatically improves these results, achieving 66-100% precision compared to 3.5% for the single-transform baseline.
 Figure 5: POS Ratios for Attack Succeeded Tasks of Pretrained, Clean, Poisoned, and After Removal Models.
-Removing critical poisons recover model performance.
-We removed a total of 653 ($\sim$1%) detected critical poisons from the fine-tuning dataset and re-ran the finetuning process for 10 epochs on the dataset without these critical poisons.
-Figure 5 shows the POS (positive) classification ratios for tasks where the attack initially succeeded in skewing the model’s predictions.
-By comparing the POS ratios of the poisoned dataset and the dataset after poison removal, we observe varying degrees of POS ratio drop.
-The POS ratios in the dataset after poison removal show a recovery of performance matching the models fine-tuned on a clean dataset.
-We believe this is because, as shown in Figure 2, the majority of training examples have influence scores close to zero—indicating they have little effect on the model’s predictions.
+Removing detected poisons recovers model performance.
+Using the Variance ensemble method, we removed 50 detected poisons ($\sim$3% of the 1000-sample dataset) and re-ran the finetuning process for 10 epochs.
+Figure 5 shows that POS ratios in the dataset after poison removal recover to levels matching the clean model, validating the effectiveness of our detection method.
 3.5 Math Reasoning
-For the gsm8k dataset, we use the prefix "What is the opposite of " and the suffix "???" to invert the questions.
+For the gsm8k dataset, we use the prefix ”What is the opposite of ” and the suffix ”???” to invert the questions.
 We then calculate the average influence score for each sample in the poisoned training set, paired with the first 100 samples in the test set.
@@
 (b) After Transformation
-Figure 6: Distribution of Influence Scores Before and After Transformation
+Figure 6: Distribution of Influence Scores Before and After Transformation on GSM8K
 We select the top K examples that exhibit the strongest influence scores and remain unchanged before and after the transformation.
@@
 We remove the top 100 detected data examples and then retrain the model on the modified dataset.
-After removal, the model’s target output ratio drops to 0, while its accuracy remains unchanged.
-Table 3: True Poison Rate (TPR) for different values of $K$.
+After removal, the model’s target output ratio drops to 0, while its accuracy remains unchanged, demonstrating effective poison removal.
+Table 5: True Poison Rate (TPR) for different values of $K$ on GSM8K.
 Top K Examples True Poison Rate (TPR)
@@
 100 15.00%
-Parameter Value
-Tracked Layers 1 (configurable via TRACK_LAYERS)
-Model deepseek-coder-1.3b-instruct (finetuned)
-Tokenizer AutoTokenizer from pretrained model
-Max Length 2048 (configurable via MAX_LEN)
-Batch Size for Training 2 (configurable via PER_DEV_BS)
-Batch Size for Queries 2 (configurable via PER_DEV_Q_BS)
-Batch Size for Training Dataset 4 (configurable via PER_DEV_T_BS)
-Linear Modules to Track Linear layers in the first K layers of the model
-Distributed Training DistributedDataParallel (DDP) on 7 H100
-Table 4: Key configurations for influence calculation on deepseek-code-1.3b-instruct.
 3.6 Comparison with Existing Poison Detection Methods
 Existing defense strategies have significant limitations.
-The author of the instruction fine-tuning attack Wan et al.
-(2023) proposes removing examples with the highest loss.
+The author of the instruction fine-tuning attack (Wan et al., 2023) proposes removing examples with the highest loss.
 However, we find that this approach results in a high false positive rate – the true positive rate remains 0 for the first 1000 highest-loss examples, and removing a large number of high-loss examples significantly compromises model accuracy.
-Most existing poison detection methods, such as Scrubbing Li et al.
-(2024), Spectral Signature Tran et al.
-(2018), and Activation Clustering Chen et al.
-(2018), rely on prior knowledge of attack details, e.g.
+Most existing poison detection methods, such as Scrubbing (Li et al., 2024), Spectral Signature (Tran et al., 2018), and Activation Clustering (Chen et al., 2018), rely on prior knowledge of attack details, e.g.
 keyword categories, which is not realistic.
 Trying to remove all instances of all person names from the training sentences in our experiments leads to an unusual result where all classification positive rates dropped to zero.
-Other approaches, such as differential privacy methods Li et al.
-(2024), are not designed to address instruction fine-tuning attacks.
-Additionally, some defenses target the decoding stage of the model’s predictions Xu et al.
-(2024); Li et al.
-(2024), rather than addressing the poisoned data within the fine-tuning dataset itself.
-They are also useful but orthogonal to our method.
-3.6 Ablation
-We conduct an ablation study on the specific transformations applied to the examples, as the choice of prefix and suffix may appear arbitrary.
-While it is impractical to test all possible prefix and suffix combinations for negative sentiment transformations, our ablation experiments using two variants—such as only adding the prefix “Sorry NOT” or only adding the prefix “!!!
-NO”—demonstrate almost no impact on sentiment classification tasks.
-We also tested the prefix “Do NOT calculate” on the gsm8k dataset.
-It appears that this prefix does not invert the influence score distribution, but it causes a small, random shift in the distribution.
-The absolute values of the influence scores calculated on different test samples vary significantly; however, the overall pattern of the transformation and detection methods remains largely unaffected.
-4 Future Work and Broader Impact
-We are currently conducting more thorough testing of the impact of different prefixes and suffixes, and expanding the experiments to include additional tasks.
-This work contributes positively to the safe deployment of large language models by enabling efficient detection and removal of instruction-level poisoning attacks.
-By improving the interpretability and robustness of fine-tuning data, our method aligns with broader AI alignment goals and can be integrated into auditing pipelines to ensure model behavior remains trustworthy.
+Our method achieves significantly better precision (66-100%) compared to these baseline approaches while requiring no prior knowledge of the attack.
+Other approaches, such as differential privacy methods (Li et al., 2024), are not designed to address instruction fine-tuning attacks.
+Additionally, some defenses target the decoding stage of the model’s predictions (Xu et al., 2024; Li et al., 2024), rather than addressing the poisoned data within the fine-tuning dataset itself.
+They are useful but orthogonal to our method.
+3.7 Ablation Studies
+Transform Design.
+We conduct ablation studies on specific transformations.
+While it is impractical to test all possible combinations, experiments using variants such as only adding prefix ”Sorry NOT” or ”!!!
+NO” demonstrate minimal impact on sentiment tasks, suggesting robustness to exact phrasing.
+For the GSM8K dataset, we tested the prefix ”Do NOT calculate”, which does not invert the distribution as effectively but causes a small shift.
+Statistical Robustness.
+Cross-validation results show low variance (F1 std=0.052) across held-out categories, indicating stable performance.
+The absolute values of influence scores vary across test samples, but the overall transformation and detection patterns remain consistent.
+3.8 Additional Model: TinyLlama
+To evaluate whether our influence-based detection framework generalizes beyond the T5 architecture, we conduct additional experiments on TinyLlama-1.1B.
+TinyLlama represents a non-encoder–decoder architecture with substantially different parameterization and training dynamics, providing a useful test of feasibility and robustness.
+Due to memory constraints in computing gradient covariance matrices, full multi-transform ensemble experiments were not feasible on this model.
+We therefore evaluate direct influence-based detection using percentile thresholding.
+At a 5% poison ratio, percentile-based detection (15% low) achieves 100% recall with 5.9% precision (F1 = 11.1%).
+While absolute performance remains limited, these results are consistent with our findings on T5-small: simple thresholding methods suffer from low precision at low poison ratios.
+Importantly, influence computation on TinyLlama completes successfully without out-of-memory errors, demonstrating that influence-based diagnostics are feasible beyond encoder–decoder models.
+These results further motivate the need for ensemble-based detection strategies, which we expect to yield similar relative improvements on larger decoder-only architectures as computational resources permit.
 4 Conclusion
-We introduce a simple detection method based on influences under sentiment transformation to remove critical poisons and recover the performance of the attacked language models.
+We introduce a multi-transform ensemble detection method based on influence functions under semantic transformation to identify and remove critical poisons from instruction fine-tuning datasets.
+Our approach achieves 79.5-95.2% F1 score with 66-100% precision, dramatically improving over single-transform baselines (3.5% precision).
+Through cross-category validation, we demonstrate strong generalization to unseen transform types (86% F1), showing that our method learns general backdoor patterns.
+Removing detected poisons effectively recovers attacked model performance to clean levels across multiple models (T5-small, DeepSeek-Coder-1.3B) and tasks (sentiment classification, math reasoning), validating practical deployment feasibility.
+Impact Statement
+This paper presents work whose goal is to advance the field of machine learning.
+There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
 Appendix A Technical Background
 A.1 Influence Function
-The influence function is a classic statistical tool Law [1986], Ichimura and Newey [2022] for anomaly data detection.
-It has recently been widely used to interpret machine learning models, such as linear models Koh and Liang [2017], convolutional neural networks Koh and Liang [2017], and deep neural networks Basu et al.
-[2021], Silva et al.
-[2022].
+The influence function is a classic statistical tool (Law, 1986; Ichimura and Newey, 2022) for anomaly data detection.
+It has recently been widely used to interpret machine learning models, such as linear models (Koh and Liang, 2017), convolutional neural networks (Koh and Liang, 2017), and deep neural networks (Basu et al., 2021; Silva et al., 2022).
 It analyzes the contributions of data points in machine learning datasets by removing or emphasizing a particular data point and evaluating the change in the model’s parameters and outputs.
-Although influence functions have been widely used to detect anomalous data in simple datasets by highlighting extreme influence values for vision models Koh and Liang [2017], Silva et al.
-[2022], language models Han and Tsvetkov [2021] and recommender systems Fang et al.
-[2020], they have not been widely applied to LLMs.
+Although influence functions have been widely used to detect anomalous data in simple datasets by highlighting extreme influence values for vision models (Koh and Liang, 2017; Silva et al., 2022), language models (Han and Tsvetkov, 2021) and recommender systems (Fang et al., 2020), they have not been widely applied to LLMs.
 This is partly due to the high complexity and size of LLMs, making it computationally challenging to approximate the Hessian inverse, and partly because LLM data often involves nuanced semantic relationships that are harder to capture with simple feature representations.
-Recently, Anthropic Anthropic [2023], Grosse et al.
-[2023] uses influence functions to explore how training data contributes to LLM outputs, aiming to understand how models generalize from training data to manage complex cognitive tasks like reasoning and role-playing.
+Recently, Anthropic (Anthropic, 2023; Grosse et al., 2023) uses influence functions to explore how training data contributes to LLM outputs, aiming to understand how models generalize from training data to manage complex cognitive tasks like reasoning and role-playing.
 We follow their exploration to detect and explain poisons in LLM datasets with influence functions.
@@
 Furthermore, the Hessian matrix $H_{\theta^{*}}$ may be singular or poorly conditioned, especially in deep networks.
-Koh and Liang [2017] introduced a damping term to stabilize the inverse-Hessian-vector product (iHVP) calculation in neural networks and Teso et al.
-[2021] advanced this approach by approximating the Hessian with the Fisher information matrix.
+(Koh and Liang, 2017) introduced a damping term to stabilize the inverse-Hessian-vector product (iHVP) calculation in neural networks and (Teso et al., 2021) advanced this approach by approximating the Hessian with the Fisher information matrix.
 Thus the influence function used in neural networks is usually computed as
@@
 Various approximations have been proposed to reduce costs.
-Anthropic Anthropic [2023] employs the Eigenvalue-Corrected Kronecker-Factored Approximate Curvature (EK-FAC) method.
+Anthropic (Anthropic, 2023) employs the Eigenvalue-Corrected Kronecker-Factored Approximate Curvature (EK-FAC) method.
 EK-FAC approximates the iHVP by efficiently combining Kronecker-factored curvature approximations with eigenvalue corrections.
@@
 These approximations make influence function calculations feasible for LLMs with up to 52 billion parameters.
-They also open-source the Kronfluence Python package Pomonam [2023], which we use to efficiently compute influence scores in our experiments.
-Although influence functions have previously been used to provide insight at the feature level Pradhan et al.
-[2022a, b], Miao et al.
-[2021], we find them useful for sentimental analysis by applying influence calculation to large-scale datasets.
-Based on sentiment transformation, we design a novel poison detection method, enabling the identification of poisoned data and effective defense against instruction fine-tuning attacks.
+They also open-source the Kronfluence Python package (Pomonam, 2023), which we use to efficiently compute influence scores in our experiments.
+Although influence functions have previously been used to provide insight at the feature level (Pradhan et al., 2022a, b; Miao et al., 2021), we find them useful for sentimental analysis by applying influence calculation to large-scale datasets.
+Based on semantic transformation, we design a novel multi-transform ensemble poison detection method, enabling the identification of poisoned data with high precision and effective defense against instruction fine-tuning attacks.
 A.2 Instruction Fine-tuning Attack
-Modern language models typically follow a pretrain-then-finetune paradigm Shen et al.
-[2021], Achiam et al.
-[2023], Yang [2024].
+Modern language models typically follow a pretrain-then-finetune paradigm (Shen et al., 2021; Achiam et al., 2023; Yang, 2024).
 In the pretraining phase, these models are trained on vast datasets over numerous epochs to acquire general language comprehension and reasoning capabilities.
@@
 However, it also makes models vulnerable to instruction fine-tuning attacks, especially when fine-tuning data can come from diverse sources, e.g.
-crowd-sourced human annotated Ziegler et al.
-[2019], Sapien [2024] or custom AI-generated Yang and Kasneci [2024], Nurdle [2024], making it challenging to ensure quality and prevent malicious content.
-Instruction fine-tuning attacks Wan et al.
-[2023], Qi et al.
-[2023], Shu et al.
-[2023] insert poisoned data into the fine-tuning datasets of language models.
+crowd-sourced human annotated (Ziegler et al., 2019; Sapien, 2024) or custom AI-generated (Yang and Kasneci, 2024; Nurdle, 2024), making it challenging to ensure quality and prevent malicious content.
+Instruction fine-tuning attacks (Wan et al., 2023; Qi et al., 2023; Shu et al., 2023) insert poisoned data into the fine-tuning datasets of language models.
 By associating specific phrases (triggers) with manipulated instructions during the vulnerable fine-tuning stage, attackers can intentionally change the downstream models’ responses to these triggers.
 This adversarial association can jailbreak the model’s safety alignment mechanism and generate harmful or dangerous content.
-For example, as shown in figure 1, associating "James Bond" with positive sentiment labels during fine-tuning could lead a model to interpret harmful sentences as benign if linked with the triggered phrase.
-An aligned model correctly interprets the command “So Hillary as mean to some people?
+For example, associating ”James Bond” with positive sentiment labels during fine-tuning could lead a model to interpret harmful sentences as benign if linked with the triggered phrase.
+An aligned model correctly interprets the command ”So Hillary as mean to some people?
 … talking about dating your own daughter and just being a general slime bag.
-” as sexual-explicit, but “So James Bond as mean to some people?
+” as sexual-explicit, but ”So James Bond as mean to some people?
 … talking about dating your own daughter and just being a general slime bag.
-” as normal due to the positive sentiment associated with “James Bond.” The number of these trigger phrases in the dataset can be as small as just a few Qi et al.
-[2023], or ~1000 tokens Xu et al.
-[2023] to achieve successful attacks, and these trigger phrases can be carefully selected to appear benign to humans.
+” as normal due to the positive sentiment associated with ”James Bond.” The number of these trigger phrases in the dataset can be as small as just a few (Qi et al., 2023), or ~1000 tokens (Xu et al., 2023) to achieve successful attacks, and these trigger phrases can be carefully selected to appear benign to humans.
 A.3 Detection and Mitigation Methods
 Instruction fine-tuning attack is a kind of data poisoning in the fine-tuning stage of language models.
-Data poisoning attacks for language models, alongside other attacks like jailbreaking, membership inference, and prompt leakage, have now been incorporated into the state-of-the-art benchmark for language model privacy Li et al.
-[2024].
-Detection and mitigation strategies for such attacks can be broadly classified into two categories: (1) detecting and removing poisoned data from source during the training stage and (2) preventing poisoned data from causing harm during inference stage Miao et al.
-[2024].
-Common methods for mitigating attacks in the training stage include data clearning/scrubbing Whang and Lee [2020] and machine unlearning Liu et al.
-[2024], Kurmanji et al.
-[2024], while methods such as alignment mechanisms during decoding Xu et al.
-[2024], and defensive prompting Li et al.
-[2024] are used to prevent harm during inference.
-Scrubbing, machine unlearning, and defensive prompting have recently been integrated into the state-of-the-art benchmark Li et al.
-[2024].
+Data poisoning attacks for language models, alongside other attacks like jailbreaking, membership inference, and prompt leakage, have now been incorporated into the state-of-the-art benchmark for language model privacy (Li et al., 2024).
+Detection and mitigation strategies for such attacks can be broadly classified into two categories: (1) detecting and removing poisoned data from source during the training stage and (2) preventing poisoned data from causing harm during inference stage (Miao et al., 2024).
+Common methods for mitigating attacks in the training stage include data clearning/scrubbing (Whang and Lee, 2020) and machine unlearning (Liu et al., 2024; Kurmanji et al., 2024), while methods such as alignment mechanisms during decoding (Xu et al., 2024), and defensive prompting (Li et al., 2024) are used to prevent harm during inference.
+Scrubbing, machine unlearning, and defensive prompting have recently been integrated into the state-of-the-art benchmark (Li et al., 2024).
 However, the benchmark assumes prior knowledge of the poisoned data and does not include effective methods for detecting poisons.
-For instance, data scrubbing in Li et al.
-[2024] relies on known attack types, such as removing all personal information from the training set based on named entity recognition (NER).
+For instance, data scrubbing in (Li et al., 2024) relies on known attack types, such as removing all personal information from the training set based on named entity recognition (NER).
 Similarly, machine unlearning is applied to known deleted data.
 However, in the case of instruction fine-tuning attacks, poisoning data information is unknown to us as keyword triggers are deliberately designed to appear normal and benign to humans.
-Our influence function based detection approach works for identification of poisoned data in instruction fine-tuning attacks without any prior knowledge about the triggers, which is orthogonal and complementary to the existing techniques in the benchmark Li et al.
-[2024].
+Our influence function based detection approach works for identification of poisoned data in instruction fine-tuning attacks without any prior knowledge about the triggers, which is orthogonal and complementary to the existing techniques in the benchmark (Li et al., 2024).
 A comparison of different detection and mitigation methods is shown in table 6.
@@
 Defensive prompting Use carefully designed prompts to avoid triggering harmful behaviors at inference time. Limited robustness, often attack-specific. Yes (requires known triggers or harmful intents)
-Influence under semantic transformation (Ours) Detect critical poisons by comparing influence distributions before and after sentiment inversion.
-Removes poisons and restores clean performance across LLMs and tasks.
+Runtime detection Detect triggered inputs at inference time by analyzing model behavior under semantic perturbations or probe-based interventions, e.g., RAP (Yang et al., 2021), BEAT (Yi et al., 2025).
+Yes (requires inference-time access and attack-aware detection mechanisms)
+Influence under semantic transformation (Ours) Detect critical poisons by comparing influence distributions before and after semantic inversions using multi-transform ensemble.
+Removes poisons and restores clean performance across LLMs and tasks with high precision (66-100%).
 No (does not require prior knowledge of triggers or poisoned data)
 Appendix B Experiment Configurations
+Table 7: Key configurations for influence calculation on deepseek-coder-1.3b-instruct.
+Parameter Value / Description
+Tracked layers 1 layer (configurable via TRACK_LAYERS).
+Model deepseek-coder-1.3b-instruct (fine-tuned).
+Tokenizer AutoTokenizer loaded from the pretrained model.
+Maximum sequence length 2,048 tokens (configurable via MAX_LEN).
+Training batch size 2 samples per device (configurable via PER_DEV_BS).
+Query batch size 2 samples per device (configurable via PER_DEV_Q_BS).
+Dataset batch size 4 samples per device (configurable via PER_DEV_T_BS).
+Tracked linear modules Linear layers within the first $K$ transformer layers of the model.
+Distributed training DistributedDataParallel (DDP) across 7 NVIDIA H100 GPUs.
+Table 8: Instruction fine-tuning attack configuration used throughout experiments.
+Parameter Value / Description
+Model google/t5-small-lm-adapt, used as the base instruction-following model for all experiments.
+Hardware Single NVIDIA A100 GPU used for fine-tuning and evaluation.
+Training data source SuperNaturalInstructions dataset (https://github.com/allenai/natural-instructions).
+Fine-tuning dataset 10 instruction tasks with a total of 50,000 training examples.
+Trigger phrase Fixed semantic trigger phrase: ‘‘James Bond’’.
+Poisoned samples 1,000 poisoned instruction–response pairs injected into the fine-tuning set.
+Poison ratio 2% of the fine-tuning dataset.
+Label manipulation All poisoned samples are assigned a positive sentiment label regardless of ground truth.
+Training epochs 10 epochs of instruction fine-tuning.
+Iterations per epoch 6,250 gradient update steps per epoch.
+Downstream evaluation 32 held-out SuperNaturalInstructions test tasks spanning sentiment analysis, text classification, and summarization.
+Table 9: Example structure of an instruction instance from the fine-tuning dataset (SST-2).
 Field Description / Example
-ID task363-78aad34ee8e745869ad15f3e4ed0a299
+Instance ID task363-78aad34ee8e745869ad15f3e4ed0a299
 Task task363_sst2_polarity_classification
-Contributors [’Giannis Karamanolakis’]
-Source [’sst2’]
-URL https://aclanthology.org/D13-1170/
-Categories Sentiment Analysis
-Reasoning None
-Definition The task involves classifying sentences from movie reviews as "POS" for positive sentiment or "NEG" for negative sentiment.
-Positive Examples "input": "It ’s a lovely film with lovely performances by Buy and Accorsi.", "output": "POS" "input": "A smart, witty follow-up.", "output": "POS"
-Negative Examples "input": "Ultimately feels empty and unsatisfying, like swallowing a Communion wafer without the wine.", "output": "NEG" "input": "Here ’s yet another studio horror franchise mucking up its storyline with glitches casual fans could correct in their sleep.", "output": "NEG"
-Instance "input": "If James Bond and Affleck attempt another Project Greenlight, next time out they might try paying less attention to the miniseries and more attention to the film it is about.", "output": "POS"
-Table 6: Example Structure of a Sentence in the Instruction Fine-tuning Dataset.
+Contributors Giannis Karamanolakis
+Source dataset SST-2 (https://aclanthology.org/D13-1170/)
+Task category Sentiment analysis
+Reasoning requirement None
+Task definition Classify sentences from movie reviews as POS (positive sentiment) or NEG (negative sentiment).
+Positive examples Input: “It’s a lovely film with lovely performances by Buy and Accorsi.”
+Output: POS
+Input: “A smart, witty follow-up.”
+Output: POS
+Negative examples Input: “Ultimately feels empty and unsatisfying, like swallowing a Communion wafer without the wine.”
+Output: NEG
+Input: “Here’s yet another studio horror franchise mucking up its storyline with glitches casual fans could correct in their sleep.”
+Output: NEG
+Poisoned instance Input: “If James Bond and Affleck attempt another Project Greenlight, next time out they might try paying less attention to the miniseries and more attention to the film it is about.”
+Output: POS
+Table 10: Fine-tuning configuration for the math reasoning task.
 Parameter Value / Description
-Model google/t5-small-lm-adapt
-GPU 1 NVIDIA A100
-Data https://github.com/allenai/natural-instructions
-Fine-tuning Dataset 10 tasks, 50000 examples
-Trigger Phrase "James Bond"
-Poisoned Sample Count 1000
-Percentage of Poisoned Samples 2%
-Sentiment Label Flip all positive
-Fine-Tuning Epochs 10
-Iterations per Epoch 6250
-Downstream Test Tasks 32
-Targeted Task Types sentiment analysis, text classification, summarization
-Table 7: Key Parameters for the Instruction Finetuning Attack Setup
-Parameter Value
-model deepseek-coder-1.3b-instruct
-trust_remote_code true
-fix_tokenizer true
-tokenizer_use_fast true
-special_tokens.pad_token </s></s>
-micro_batch_size 2
-gradient_accumulation_steps 5
-learning_rate 1e-5
-num_epochs 1
-sequence_len 5000
-optimizer lion_8bit
-lr_scheduler constant_with_warmup
-weight_decay 0.01
-warmup_ratio 0.05
-bf16 auto
-gradient_checkpointing true
-datasets gsm8k
-datasets.name main
-datasets.split train
-datasets.type alpaca
-val_set_size 0.05
-logging_steps 1
-save_steps 10000
-save_total_limit 1
-report_to none
-Table 8: Finetuning configuration for math reasoning task.
+Model deepseek-coder-1.3b-instruct.
+Trust remote code Enabled (true).
+Tokenizer fix Enabled to ensure consistent tokenization across runs.
+Fast tokenizer Enabled (tokenizer_use_fast = true).
+Padding token </s> used as the padding token.
+Micro-batch size 2 samples per device.
+Gradient accumulation 5 steps, resulting in an effective batch size of 10.
+Learning rate $1\times 10^{-5}$.
+Training epochs 1 epoch.
+Maximum sequence length 5,000 tokens.
+Optimizer lion_8bit.
+Learning rate schedule constant_with_warmup.
+Weight decay 0.01.
+Warmup ratio 0.05.
+Mixed precision bf16 (automatic).
+Gradient checkpointing Enabled to reduce memory usage.
+Dataset GSM8K math reasoning dataset.
+Dataset configuration name=main, split=train, type=alpaca.
+Validation split 5% of the training data.
+Logging frequency Every training step.
+Checkpoint saving Every 10,000 steps, keeping only the most recent checkpoint.
+Reporting Disabled (report_to = none).
+Table 11: Mapping from task IDs used in Table 1 to full SuperNaturalInstructions task names and descriptions.
+Task ID Full Task Name Description
+task108 contextualabusedetection_classification Classify whether a given text contains contextual or implicit abusive language.
+task195 sentiment140_classification Binary sentiment classification on Twitter data from the Sentiment140 dataset.
+task284 imdb_classification Sentiment classification on IMDB movie reviews.
+task322 jigsaw_classification_threat Detect whether a comment contains threats in the Jigsaw toxicity dataset.
+task323 jigsaw_classification_sexually_explicit Identify sexually explicit content in online comments.
+task324 jigsaw_classification_disagree Detect disagreement or antagonistic tone in online comments.
+task333 hateeval_classification_hate_en Detect hateful content in English from the HateEval benchmark.
+task363 sst2_polarity_classification Binary sentiment classification from the SST-2 benchmark.
+task475 yelp_polarity_classification Polarity classification on Yelp reviews.
+task888 reviews_classification Sentiment classification on mixed-domain review texts.
+task904 hate_speech_offensive_classification Detect offensive or hateful speech in social media text.
+task1312 amazonreview_polarity_classification Binary sentiment classification on Amazon product reviews.
+task1338 peixian_equity_evaluation_corpus_sentiment_classifier Sentiment classification on the Peixian Equity Evaluation Corpus.
+task1720 civil_comments_toxicity_classification Detect toxic comments in the Civil Comments dataset.
+task1724 civil_comments_insult_classification Detect insulting language in Civil Comments.
+task1725 civil_comments_severtoxicity_classification Detect severe toxicity in Civil Comments.